### v1

# Generalized Bayesian deep reinforcement learning

Abstract
Bayesian reinforcement learning (BRL) is a method that merges principles from Bayesian statistics and reinforcement learning to make optimal decisions in uncertain environments. Similar to other model-based RL approaches, it involves two key components: (1) Inferring the posterior distribution of the data generating process (DGP) modeling the true environment and (2) policy learning using the learned posterior. We propose to model the dynamics of the unknown environment through deep generative models assuming Markov dependence. In absence of likelihood functions for these models we train them by learning a generalized predictive-sequential (or prequential) scoring rule (SR) posterior. We use sequential Monte Carlo (SMC) samplers to draw samples from this generalized Bayesian posterior distribution. In conjunction, to achieve scalability in the high dimensional parameter space of the neural networks, we use the gradient based Markov chain Monte Carlo (MCMC) kernels within SMC. To justify the use of the prequential scoring rule posterior we prove a Bernstein-von Misses type theorem. For policy learning, we propose expected Thompson sampling (ETS) to learn the optimal policy by maximising the expected value function with respect to the posterior distribution. This improves upon traditional Thompson sampling (TS) and its extensions which utilize only one sample drawn from the posterior distribution. This improvement is studied both theoretically and using simulation studies assuming discrete action and state-space. Finally we successfully extend our setup for a challenging problem with continuous action space without theoretical guarantees.
1 Introduction
Effective learning and decision-making within perpetually changing dynamic systems are imperative for applications like controlling automated machinery and enabling robotic navigation. Reinforcement learning (RL) stands out as a potent tool in these domains, enabling agents to acquire knowledge through trial and error and responses from the environment. Its versatility has resulted in its widespread utilization across a range of sectors, encompassing automated vehicles (Guan et al., 2020), robotics (Kormushev et al., 2013), healthcare (Yu et al., 2021), finance (Deng et al., 2016), and various natural language processing (NLP) applications (Uc-Cetina et al., 2023), recommender systems (Chen et al., 2023) and so on.
Typically, an RL task can be expressed using a Markov decision process (MDP) (Sutton and Barto, 2018) framework. MDPs are a classical framework for sequential decision-making, where a policy maps the state space to the action space. The multitude of RL algorithms existing in the literature can be broadly segregated into two categories: those reliant on a model (‘model-based’ algorithms) and those that facilitate policy learning in a ‘model-free’ manner. Model-free algorithms like Q-learning algorithms (Clifton and Laber, 2020) and policy gradient methods (Sutton and Barto, 2018) directly learn from the history generated by real time interactions. Conversely, model-based approaches like dynamic programming (Sutton and Barto, 2018), Monte Carlo tree search (Coulom, 2006), PILCO (Deisenroth and Rasmussen, 2011) etc. use knowledge about the environment to design a policy.
When an efficient model of the environment is available or can be learned, model-based RL (Moerland et al., 2023) strategies are more sample-efficient than model-free methods. In practical scenarios, explicit models for environment dynamics are often unavailable, but simulator models may be accessible. In the recent advances of model-based RL, neural network based predictive models (Nagabandi et al., 2018), (Kaiser et al., 2019); conditional GANs (Charlesworth and Montana, 2020), (Zhao et al., 2021) have been successfully adopted to model the complex dynamical systems underlying the environment. These models are extremely potent as seen in a multitude of recent applications, but their complexity makes training challenging due to the lack of a tractable objective function. Hence previous work either approximates a divergence function (as in adversarial training regime (Goodfellow et al., 2020) or the Fisher score using a function class parametrized by neural networks (Gurney, 2018). Both of these strategies are sample-greedy and computationally expensive defeating the purpose of online RL, further the adversarial training is known to be notoriously difficult to train and faces issues (e.g. mode collapse (Arora et al., 2018)) due to the accumulation of the bias in each iteration of training.
For more robust training of the underlying model, probabilistic model predictive control (MPC) (Kamthe and Deisenroth, 2018), conditional variational inference (VI) to capture multimodal stochasticity (Moerland et al., 2017), and Bayesian neural networks (BNNs) (Depeweg et al., 2016) etc have been explored in the literature. Notably, the Bayesian approach offers the additional benefit of quantifying uncertainty in the model’s learning of the environment’s dynamics. However, in real-world scenarios, the likelihood function required for these updates is often intractable, necessitating likelihood-free inference.
In this work we mainly focus on Bayesian model-based RL approach which integrates Bayesian inference with RL to make decisions in uncertain environments. It comprises two primary components: (1) Bayesian inference of the underlying MDP that governs environment dynamics, and (2) designing a policy based on the learned model. Hence, our main contribution is in providing a likelihood-free framework for training Bayesian generative NN to model the underlying MDP and a strategy to use this model for policy learning.
Bayesian inference of the MDP
To perform likelihood-free inference, we define a generalized Bayesian scoring rule posterior (Pacchiardi et al., 2024b) using predictive-sequential (or prequential) scoring rules (Dawid, 1984; Dawid and Vovk, 1999; Dawid, 1991). Prequential SRs evaluate the fit of observed data to predictions conditioned on the past data using strictly proper scoring rules (e.g. energy score, kernel score etc., see (Gneiting and Raftery, 2007) for a full review). Prequential SR has been previously used to train generative network based models for probabilistic forecasting (Pacchiardi et al., 2024a). The same idea can be extended to define a generalized posterior which uses the prequential SR as an alternative for the negative log likelihood of a model. We further show that a Bernstein-von Mises type theorem (showing contraction of the posterior to the true parameter value) holds for these generalized posteriors when a deep generative MDP is used to model the environment. Even in the absence of a tractable likelihood function, an unbiased estimate of the SR can be derived using the observed data and samples from the generative models. Further, when the generative model is differentiable, an unbiased estimator of the gradient of the generalized posterior is easy to derive which enables the exploration of the parameter space in an efficient and scalable manner.
Markov Chain Monte Carlo (MCMC) methods are classical tools for sampling from a posterior distribution. Since the observed data in reinforcement learning (RL) scenarios is typically acquired sequentially through interactions with the environment, we opt for SMC samplers (Del Moral et al., 2006) to infer the posterior distribution of model parameters. To enhance the performance of SMC samplers, particularly in the high-dimensional parameter spaces of the generative networks, we incorporate gradient-based MCMC kernels to move the particles. By computing an unbiased estimator of the gradient of the prequential scoring rule, we can leverage the extensive literature on stochastic gradient MCMC methods (Nemeth and Fearnhead, 2021). Specifically, we utilize preconditioned adapted stochastic gradient Langevin dynamics (SGLD) (Jones and Leimkuhler, 2011; Patterson and Teh, 2013; Chen et al., 2016), which are well-suited for exploring high-dimensional parameter spaces. However, when the simulator model is non-differentiable, a common challenge in many RL problems, gradient-based kernels become problematic. In such cases, we can employ zeroth-order gradient estimates (Liu et al., 2020), which rely solely on simulations from the generative models, though this may introduce some estimation bias and variance.
Policy search
Bayesian RL using Thompson sampling (Russo et al., 2018) and their extensions (Osband et al., 2013) are known to have excellent performance with good regret bounds, but while training the policy to explore the parameter space they only utilize one sample drawn from the posterior distribution. However, there is sparse evidence in the literature (Dimitrakakis and Ortner, 2022) showing improved performance whilst using more samples from the inferred posterior distribution. Building upon this, we propose a new policy learning methodology called expected Thompson sampling (ETS) which learns the optimal policy by maximizing the expected value function w.r.t. the inferred posterior distribution. In practice, we estimate the expected value function using samples drawn from the posterior. We also show both using simulation based and theoretical studies that the regret of the policy learned via ETS decreases as the number of posterior samples increases.
Therefore, our proposed approach combines the robustness of scoring rules for Bayesian inference, employs gradient-based SMC for high dimensional parameter spaces, and provides a strategy to design a policy, by potentially better handling model uncertainty. The main contributions of our work are as follows:
  • A likelihood-free framework is proposed to enable the ETS for MDPs without likelihoods (eg. deep generative models) and we show that it is more sample-efficient compared to model-free methods.
  • We provide a theoretical bound for the approximation error in the action value function due to using ETS for generalized posterior.
  • While ETS-based policy learning is computationally intensive due to posterior sampling of model parameters, we make it feasible and scalable by utilizing the SMC for the sequential learning task in hand and gradient based MCMC kernels to handle high-dimensional parameter space.
Figure 1: Bayesian deep RL: The above diagram shows an episodic posterior update followed by a policy update routine for $k$th episode ($k=1,2,\ldots$). Let us assume the episode length to be $\tau$, and $\pi_{0}({\theta})$ to be the prior on the model parameters. Prequential Scoring Rule ($\mathcal{S}$) is used to compute the generalized posterior which is based on true interaction data(${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X}$) and a simulation(${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X(\theta)}$) from the deep generative model $\mathcal{M}(\theta)$. We obtain the samples, $\{\theta_{1},\theta_{2},\ldots,\theta_{n}\}$ from the generalized posterior ($\pi_{k}(\theta)\propto\pi_{k-1}(\theta)\exp(-\mathcal{S}(X^{(k)},X(\theta)))$) via Sequential Monte Carlo (SMC) samplers. These samples are used to simulate $n$ trajectories of interaction (${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X(\theta_{1}),X(\theta_{2}),\ldots,X(\theta_{n})}$) from the model $\mathcal{M}(\theta)$. Optimal policy is trained by maximizing the averaged action value function computed from the $n$ simulated trajectories. The new policy, $\mu^{(k+1)}$ is then used to interact with the true Environment in the next episode.
In Section 2 we provide an overview of Bayesian model-based RL and the relevance of Thompson sampling. In Section 2.1 we introduce expected Thompson sampling (ETS) and demonstrate its application on a simple “chain-task” example for which the likelihood function is available. Section 3 details the generalized Bayesian inference framework, with Section 3.2 discussing the properties of the prequential scoring rule posterior and Section 3.3 explaining the SMC sampler used to sample from the posterior. In Section 4, we derive an error bound for the approximate action-value function under the ETS framework. Simulation studies are presented in Section 5 to evaluate the approach, and Section 6 summarizes key findings and conclusions.
2 Bayesian Model-based Reinforcement learning
In RL, an agent (eg. a video game player) interacts with an unknown environment (eg. the virtual reality inside the game) by taking actions causing the transition of the environment to a new state, yielding rewards in the process. Suppose, at time $t$, the environment was observed at the state $s_{t}\in\mathcal{S}$, where $\mathcal{S}$ is the state space. After the agent’s action $a_{t}\in\mathcal{A}$ (the action space $\mathcal{A}$), the environment moves to the next state, $s_{t+1}\in\mathcal{S}$, and the agent receives a reward $r_{t+1}\in\mathcal{R}$ where $\mathcal{R}$ is the reward space. The goal of the agent is to devise a strategy that selects actions based on the current state to maximize cumulative rewards. Under the assumption that the new state ($s_{t+1}$) only depends upon the previous state and action ($s_{t},a_{t}$) and not on $\{s_{u},a_{u}:u<t\}$ (Markovian assumption), the environment can be considered as a Markov decision process (MDP), as defined below.
A Markov decision process (MDP) $\mathbb{M}$ on the state space $\mathcal{S}$ and action space $\mathcal{A}$, can be defined by the distribution of the initial state $\rho$ and the transition probabilities $P(S_{t+1}=s_{t+1}|S_{t}=s_{t},A_{t}=a_{t})$ [^1] Here, uppercase letters denote random variables, with their lowercase versions representing their realizations. and $P(R_{t+1}=r_{t+1}|S_{t}=s_{t},A_{t}=a_{t})$. The ultimate objective is to develop a policy $\mu$ that maximizes the total reward over the long term in the future, where $\mu:\mathcal{S}\rightarrow\mathcal{A}$ is a mapping of the state space to the the action space. This involves making decisions—choosing actions based on the current state—to navigate the environment in a way that accrues the maximum possible reward. Consequently, it becomes an optimization problem where the objective function is dependent on future rewards. To formalize this, we derive the value of a policy $\mu$ as the expected discounted return from any initial state if the policy $\mu$ is followed thereafter to interact with the environment. Hence, for any time $t$, we define the discounted return as $G_{t}=\sum_{k=0}^{\infty}\gamma^{k}{R}_{t+k+1}$. Then the value function of a policy $\mu$ at the state $S_{t}=s$, is defined as,
$V^{\mathbb{M}}_{\mu}(s)=\mathbb{E}_{\mathbb{M},\mu}\left[G_{t}|S_{t}=s\right]\quad\text{for all }s\in\mathcal{S}$
Here $\gamma\in(0,1]$ is a discounting factor. A policy $\mu$ is said to be optimal for $\mathbb{M}$ if, $V_{\mu}^{\mathbb{M}}(s)=$ $\underset{\mu^{\prime}}{\operatorname{max}}V_{\mu^{\prime}}^{\mathbb{M}}(s)$ for all $s\in\mathcal{S}$.
However, as is often the case, instead of directly calculating the value function of a policy, we calculate the action-value function. For any time $t$, the action-value function of a policy $\mu$ for the state-action pair, $(S_{t}=s,A_{t}=a)$ is defined as,
$Q^{\mathbb{M}}_{\mu}(s,a)=\mathbb{E}_{\mathbb{M},\mu}\left[G_{t}|S_{t}=s,A_{t}=a\right]\quad\text{for all }s\in\mathcal{S},a\in\mathcal{A}$
A natural way to assess the performance of a policy is via its regret. Regret quantifies the performance gap between a learned policy and the optimal policy, showing how much is lost due to suboptimal decisions. Let us denote the optimal policy under the MDP $\mathbb{M}$ by $\mu_{0}$. Then, the regret of a policy $\mu$ is given by, $\text{Regret}(\mu)=\sum_{s\in\mathcal{S}}\rho(s)(V^{\mathbb{M}}_{\mu_{0}}(s)-V^{\mathbb{M}}_{\mu}(s))$
Parametric MDP
Calculation of the value function is not straightforward since it involves taking expectations with respect to the MDP, which is not known in advance. Common approaches include using Bellman equations to approximate the value function based on observations from the environment received through interactions. These approaches are often sample-inefficient, as precise approximation of expectations requires extensive training data. To mitigate this sample inefficiency, in model based RL a parametric model for the MDP is assumed. If we can learn a good model emulating the true MDP, this can solve the above issue as we can draw many samples from the parametric MDP from which the optimal policy can be learned. But this efficiency can only be achieved if we can learn a good model even with smaller training datasets, which we answer positively here. A parametric MDP $\mathbb{M}_{\theta}$, parameterized by $\theta=(\theta_{st},\theta_{r})\in\Theta$, is defined over the state space, $\mathcal{S}$ and action space $\mathcal{A}$. We further assume the action space being finite, meaning $\mathcal{A}$ is countable and $|\mathcal{A}|=n_{a}<\infty$. [^2] Our proposed method and theoretical work will be developed under this assumption, but we will show empirically an extension of this for continuous action space, leaving the theoretical development as future work. $\mathbb{M}_{\theta}$ can be described by the distribution of the initial state $\rho$ and
$S_{t+1}|S_{t}=s_{t},A_{t}=a_{t}$ $\sim P_{\theta_{st}}(.|s_{t},a_{t})$
$R_{t+1}|S_{t}=s_{t},A_{t}=a_{t}$ $\sim P_{\theta_{r}}(.|s_{t},a_{t})$
The distributions $P_{\theta_{st}}$ and $P_{\theta_{r}}$ parameterized by $\theta_{st}$ and $\theta_{r}$ model the state transitions and reward distributions, respectively, for $\mathbb{M}_{\theta}$. For simplicity, the value function and the action-value function of a policy $\mu$ under $\mathbb{M}_{\theta}$ will be denoted as $V^{\theta}_{\mu}$ and $Q^{\theta}_{\mu}$ respectively . Let =$\mathcal{M}=\{\mathbb{M}_{\theta}:\theta\in\Theta\}$ represent a class of Markov decision processes parameterized by $\theta$. If the true underlying MDP, denoted by $\mathbb{M}_{0}$, does or does not belong to the model class $\mathcal{M}$, then we call $\mathbb{M}_{0}$ being well-specified or misspecified respectively.
Bayesian inference of parametric MDP
Assume we have generated a trajectory of state-action-reward-next state sequences of length $T$ through interacting with the unknown environment $\mathbb{M}$, ${H}_{T}=\{(s_{t},a_{t},r_{t+1},s_{t+1})\}_{t=1}^{T}$. To learn $\mathbb{M}_{\theta}$ or the value of the parameters $\theta$ from ${H}_{T}$, we would first define the log likelihood function for $\theta$ with respect to the observed trajectory under $\mathbb{M}_{\theta}$ as,
$L_{T}(\theta)=\sum_{t=1}^{T}\log P_{\theta_{st}}(s_{t+1}|s_{t},a_{t})+\log P_{\theta_{r}}(r_{t+1}|s_{t},a_{t}).$ (1)
Following Bayesian inference, we would update prior beliefs about the model parameters with the likelihood function using Bayes theorem
$\log\pi(\theta|{H}_{T})\propto\log\pi(\theta)+L_{T}(\theta)$ (2)
where the prior distribution $\pi(\theta)$ quantifies our prior belief and posterior distribution $\pi(\theta|{H}_{T})$ is our updated belief.
Bayesian model-based RL uses the posterior distribution of parametric MDP $\mathbb{M}_{\theta}$ learnt via Bayesian inference, to learn an optimal policy under the unknown uncertain environments. We first introduce Thompson sampling, a popular approach which only uses one parametric MDP sampled from the posterior distribution, and then propose expected Thompson sampling in Section 2.1 which instead uses many independent and identically drawn parametric MDPs from the posterior distribution.
Thompson sampling (TS)
Thompson sampling (TS)(Thompson, 1933) is a widely used approach in multi-armed bandit (MAB) problems (Lattimore and Szepesvári, 2020). In a typical MAB setup, each arm (or action) is associated with a reward, drawn from an unknown distribution. The agent’s objective is to maximize the accumulated reward through interactions with this environment. To gain insight into the reward distribution, the agent must experiment by trying different arms.
A Bayesian solution to this challenge begins with a prior belief on the mean rewards, representing our initial knowledge about the environment. As the agent interacts with the environment, it collects a sequence of action-reward pairs. This information is used to update the prior, refining the posterior distribution of the rewards. The posterior distribution is used to design a policy based on this updated knowledge.
In TS, the agent draws a sample from the updated posterior of the reward distribution and chooses the arm associated with the highest sampled reward to interact with the environment in the next round. Thus, TS offers an effective method address the exploration-exploitation dilemma in MAB problems. This approach has been generalized to reinforcement learning (RL) tasks as well (Gopalan and Mannor, 2015; Ouyang et al., 2017). In RL contexts, alongside rewards, the agent observes the environment’s state. In this case, the posterior distribution of the MDP parameters, $\theta$ becomes the quantity of interest.
2.1 Expected Thompson sampling
Thompson Sampling (TS) balances exploration and exploitation by updating a policy based on a sample from the posterior distribution of the model parameters. However, relying on a single posterior sample may introduce excessive noise, making it difficult to fully exploit the information within the posterior. A more stable and reliable approach would be to estimate the value function using multiple posterior samples. As the number of samples increases, the standard error decreases, leading to a more accurate estimate. Building on this idea, we introduce the expected Thompson sampling (ETS) algorithm, an extension of TS that leverages multiple samples from the posterior to provide a more reliable estimate of the underlying value function.
While updating the posterior at each interaction round can be computationally expensive, we mitigate this by updating the posterior after a set amount of agent-environment interaction. This makes the learning process more efficient without sacrificing performance.
In reinforcement learning (RL) tasks, terminal states often signify the end of an episode, which restarts from an initial state. These episodes, each beginning at an initial state and ending at a terminal state, are mutually independent. Tasks with clear terminal conditions, such as games that end with a win or loss, are referred to as episodic tasks. In such tasks, it is standard to update the policy at the end of each episode and use the updated policy in future interactions.
In contrast, infinite-horizon tasks have no defined terminal state. Here, a custom interaction length can be chosen to determine when to update the policy. Whether the task is episodic or not, we treat each interaction sequence as an episode and, for simplicity, assume a constant episode length $\tau$ in our analysis. However, our theorems can easily be extended to handle varying episode lengths. Next, we describe our episodic policy update strategy using ETS.
In settings where the action space $\mathcal{A}$ is finite and discrete, the optimal policy is found by optimizing the $Q$-function. In model-free approaches, data from real interactions with the environment is used for $Q$-function estimation. However, in Thompson-sampling-based methods (Dimitrakakis and Tziortziotis, 2013), a sample is drawn from the posterior of the model parameters to simulate longer trajectories (which means more data) of interactions. The $Q$-function is then estimated based on these simulated trajectories.
In the ETS framework, we extend this idea by drawing multiple samples from the posterior distribution. These posterior samples are used to simulate multiple interaction trajectories, leading to a more accurate estimate of the $Q$-function. If the posterior distribution of the model parameters after the $k$th episode is $\pi_{k}(\theta)=\pi(\theta|H_{\tau k})$, and $\bm{\theta}^{(k)}=\{\theta^{ki}\}_{i=1}^{n}$ represents a set of $n$ posterior samples, then the $Q$-function for policy $\mu$, based on the simulated trajectories, is estimated as:
[…]
Policy iteration
Policy iteration methods are a widely used class of algorithms for solving problems with discrete state spaces. These methods begin with an initial policy—often chosen at random—and then evaluate its performance by calculating the corresponding Q-function, a process known as the policy evaluation step. Following this, a new and improved policy is generated by selecting actions that maximize the estimated Q-function; this is referred to as policy improvement. The agent alternates between these two steps, iterating until the policy converges, at which point the algorithm reaches a solution.
ETS can be easily integrated with policy iteration methods by performing the policy evaluation using equation (3). In the following, we provide pseudocode for a policy iteration method incorporating ETS.
Algorithm 1 Policy iteration using expected Thompson sampling
Input: Prior distribution, $\pi$
Generate $H_{\tau}=\{(s_{t},a_{t},r_{t+1},s_{t+1})\}_{t=1}^{\tau}$ via random policy ($=\mu^{(1)}$, say )
Update the posterior, $\log\pi_{1}(\theta)\propto\log\pi(\theta)+L_{\tau}(\theta)$
for episodes $k=1,2,\ldots$ do
     Sample $\bm{\theta}^{(k)}=\{\theta^{ki}\}_{i=1}^{n}\sim\pi_{k}(\cdot)$
     Consider the initial policy $\mu_{0}=\mu^{(k)}$
     for $j=1,2,\ldots J$ do
         Compute $Q^{\bm{\theta}^{(k)}}_{\mu_{j-1}}=\frac{1}{n}\sum_{i=1}^{n}Q^{{\theta}_{ki}}_{\mu_{j-1}}$
         Update the policy $\mu_{j}(s)=\arg\max_{a}Q^{\bm{\theta}^{(k)}}_{\mu_{j-1}}(s,a)$
     end for
     Set the policy for next episode: $\mu^{(k+1)}=\mu_{J}$
     for timesteps $t=1,2,\ldots,\tau$ do
         Play $a_{\tau(k+1)+t}=\mu^{(k+1)}(s_{\tau(k+1)+t})$
         Observe $r_{\tau(k+1)+t+1}$ and $s_{\tau(k+1)+t+1}$
     end for
end for
Example
To demonstrate the effectiveness of the expected Thompson Sampling algorithm, here we use the ‘chain task’ from Dimitrakakis and Ortner (2022). The task has two actions and five states, as shown in Fig.(2). The task always starts from the leftmost state ($s(1)$, say) where the mean reward is $0.2$. There are no rewards assigned to the intermediate states. The mean reward at the terminal state (rightmost state) is 1. The first action which is denoted by the dashed-blue line takes the agent to the right, whilst the other action denoted by the red-solid line takes the agent to the first state. However, there is a probability $0.2$ that actions act in a reverse way in the environment.
0.20.20.20.20000001111
Figure 2: The chain task (Adapted from Figure 7.3 in Dimitrakakis and Ortner (2022))
The chain task is a very simple task which has the typical exploration-exploitation dilemma. Although there is a small reward at the first state, there is a bigger reward for reaching the last state. For a fairly big horizon, taking the right action is the optimal policy which we want our algorithm to learn.
Note that, here, the true model for the environment’s dynamics can be expressed with $Binomial(p_{0}=0.8)$ transition probability and rewards distributed as $\mathcal{N}(\overline{r}_{0},I_{5})$, with $\overline{r}_{0}=(0.2,0,0,0,1)$ for the 5 states. Here we assume the parametric model ($\mathbb{M}_{\theta}$) for the underlying MDP is known to us except for the true model parameters, $\theta_{0}=(p_{0},\overline{r}_{0})$. Since the true underlying MDP $\mathbb{M}_{\theta_{0}}$ belongs to the class $\mathbb{M}_{\theta}$, this is an example of well-specified model. To infer the model parameters, we started with assigning a Beta prior on the transition probabilities and a Gaussian prior on the mean rewards. Using the conjugacy of the prior distributions, we drew samples from the exact conjugate posterior distribution to perform ETS.
For policy learning, we have used a dynamic programming algorithm called Backwards Induction (BI) following the implementation of Dimitrakakis and Ortner (2022). This method is well-suited for a finite-horizon RL task with finite state and action space. The BI algorithm uses Bellman’s equations to derive the action value function starting from the terminal state and going backwards to calculate it for each state recursively. To accommodate ETS, within each episode of interaction, for each state, we calculate the $Q$ function based on the all the sampled MDPs according to equation (3) and choose the greedy action that maximizes the pooled estimate of the $Q$ function.
Now let $\mu_{0}$ be the optimal policy under the true MDP which is unknown to the agent. Suppose the agent interacts with the environment over epochs or episodes, $k=1,2,\ldots$ and $\hat{\mu}_{k}$ is the policy to determine action from a given state during the $k$th episode using ETS. The total regret of ETS due to using the sequence of policies $\hat{\mu_{1}},\hat{\mu_{2}},\dots$ over a total of $T$ timesteps is defined as the sum of episodic regrets as shown below,
[…]
$=\sum_{k=1}^{[T/\tau]}\Delta_{k,s(1)},$
since an episode always starts from the leftmost state, $s(1)$.
(a) Plot of total regret for different posterior sample size
(b) Plot of episodic regrets over the episodes
Figure 3: The chain experiment was conducted for $K=100$ episodes with an episode length of $\tau=20$. The same experiment was repeated 30 times, with the average values plotted in these graphs.
Fig. (3(a)) illustrates that total regret decreases as we increase the number of samples from the posterior to evaluate the action value functions to learn the optimal policy. Similarly, in Fig. (3(b)), we observe that the sequence of episodic regrets $\{\Delta_{k,s(1)}\}$ vanishes more rapidly as we incorporate more samples to estimate the value function, thus reducing the estimation variability that supports the statement of Theorem 3.
Hence, these figures demonstrate that learning the optimal policy can be accelerated by taking larger samples from the posterior to estimate the value functions. Consequently, the ETS algorithm outperforms standard Thompson Sampling (TS) in terms of lower regret. The ETS method operates on the principle that utilizing multiple samples from the posterior can yield a more stable approximation of the value function. Despite this advantage, traditional TS has been more widely adopted because of its lower implementation cost, even though it may suffer from approximation errors.
3 (Generalized) Bayesian model-based RL
Bayesian model-based RL requires a Bayesian treatment of the model parameters. However, when dealing with complex models with intractable likelihoods (e.g. when the analytical form for $P_{\theta_{st}}$ and $P_{\theta_{r}}$ is unavailable for a parametric MDP), directly calculating the posterior distribution becomes challenging. Nevertheless, for most of the assumed parametric MDP models we can directly sample from them, a distinctive feature that we capitalize upon. This capability to derive samples directly from the model suggests using a likelihood-free inference framework. This approach entails generating data by running simulations based on the model parameters. The samples we generate from the simulator contain a spectrum of potential outcomes under the model, given the population parameters. By using methods available under the broad umbrella of likelihood-free inference (LFI), these samples let us make inferences without having to confront the mathematical intractability of the likelihood function. Two of the most popular classes of LFI methods are approximate Bayesian computation (ABC) (Lintusaari et al., 2017) and Bayesian synthetic likelihood (BSL) (Price et al., 2018), of which ABC has been used before for Thompson Sampling in (Dimitrakakis and Tziortziotis, 2013). They approximate the intractable likelihood function either implicitly or explicitly using these samples. The asymptotic contraction of these approximate posteriors towards the true parameter value depends upon the choice of the summary statistics and some conditions being satisfied by the chosen summary statistics (Frazier et al., 2018; Li and Fearnhead, 2018; Frazier et al., 2023), which are difficult to verify in practice.
In this work, we choose a different likelihood approximation method, the scoring rule posterior framework of Pacchiardi et al. (2024b) which facilitates likelihood-free inference using a suitable scoring rule for a given type of data. This is similar to BSL in the sense that it provides an explicit approximation to the likelihood but it has outperformed both BSL and ABC regarding computational efficiency specifically more for high-dimensional examples. Further, it is easy to verify that the scoring rule posterior contracts to the true parameter value when a strictly proper scoring rule is chosen (Pacchiardi et al., 2024b). Next, we explain the scoring rule posterior and extend them for the inference of parametric MDPs, by using scoring rule posteriors based on prequential scoring rules as introduced in Pacchiardi et al. (2024a).
3.1 Scoring rule posterior
In the setup of LFI, as the likelihood function $p(\mathbf{y}^{obs}\mid\theta)$ for a model with observed data $\mathbf{y}^{obs}$ is intractable, we are unable to obtain the posterior distribution using Bayes Theorem
$\pi(\theta\mid\mathbf{y}^{obs})\propto\pi(\theta)p(\mathbf{y}^{obs}\mid\theta)=\pi(\theta)\exp\left\{\log p(\mathbf{y}^{obs}\mid\theta)\right\},$
when parameters $\theta$ and $\pi(\theta)$ are the parameters and prior distribution on our parameters. To mitigate this, Pacchiardi et al. (2024b) consider loss functions $S(P,\mathbf{y})$ known as scoring rules (Gneiting and Raftery, 2007) that measure the fit between the distribution $P_{\theta}$ of the data under parameter $\theta$, and an observed data point $\mathbf{y}$.
The scoring rule posterior is then defined as follows:
$\pi_{S}\left(\theta\mid\mathbf{y}^{obs}\right)\propto\pi(\theta)\exp\left\{-wS\left(P_{\theta},\mathbf{y}^{obs}\right)\right\}$
Comparing the two expressions, we note that the (negative) log-likelihood function can itself be considered a scoring rule (known as log-score, Dawid and Musio, 2014) and that we have introduced an additional parameter $w$, which is known as the learning rate in generalized Bayesian inference (Holmes and Walker, 2017) controlling the relative weighting of the observations relative to the prior. For this study we keep $w$ fixed at $1$. However, $w$ could be further tuned to adapt the contraction rate of the scoring rule posterior (Bissiri et al., 2016).
A scoring rule is said to be proper if its expected value with respect to the observed data $\mathbf{y}^{obs}$ is minimized when the assumed distribution $P_{\theta}$ is equal to the distribution $P_{0}$ generating the observed data $\mathbf{y}^{obs}$. If $P_{\theta}=P_{0}$ is the unique minimum, the scoring rule is said to be strictly proper. A strictly proper scoring rule is desired as this property further ensures that the best prediction one can make, according to the scoring rule, is exactly the truth. The existence of a unique minimum, i.e., the true distribution, provides a strong motivation for models to accurately capture the underlying distribution of the data. Theoretical properties of asymptotic normality and generalization bound for scoring rule posteriors when the scoring rule is strictly proper have been studied in Giummolè et al. (2019); Pacchiardi et al. (2024b). Additionally, the scoring rule posterior for some scoring rules (e.g. energy score or kernel score) exhibits robustness against outliers compared to the standard Bayes posterior using log-score.
Some examples of scoring rules are the Continuous Ranked Probability Score (CRPS) (Székely and Rizzo, 2005), Energy score or Kernel score (Gneiting and Raftery, 2007) etc. For later sections, we will be using the energy score, which can be seen as a multivariate generalization of CRPS, defined as,
$S_{E}^{(\beta)}(P_{\theta},\mathbf{y})=2\cdot\mathbb{E}||X-\mathbf{y}||_{2}^{\beta}-\mathbb{E}||X-X^{\prime}||_{2}^{\beta},\quad X,X^{\prime}\sim P_{\theta};\quad\beta\in(0,2)$ (4)
This is a strictly proper scoring rule for the class of probability measures $\mathcal{P_{\theta}}$ for which $\mathbb{E}_{X\sim P_{\theta}}||X||^{\beta}<\infty$ for all $P_{\theta}\in\mathcal{P}$ (Gneiting and Raftery, 2007). The related divergence is the square of the energy distance, which is a metric between probability distributions. Further the energy score can be unbiasedly estimated using $\mathbf{x}_{j}\sim P,j=1,\ldots,m$ which are independent and identically distributed samples from the model $P$. The unbiased estimate for the energy score can be obtained by Monte Carlo estimates of the expectations in $S_{\mathrm{E}}^{(\beta)}(P,y)$
$\hat{S}_{\mathrm{E}}^{(\beta)}\left(\left\{x_{j}\right\}_{j=1}^{m},\mathbf{y}\right)=\frac{2}{m}\sum_{j=1}^{m}\left\|x_{j}-\mathbf{y}\right\|_{2}^{\beta}-\frac{1}{m(m-1)}\sum_{\begin{subarray}{c}j,k=1\\ k\neq j\end{subarray}}^{m}\left\|x_{j}-x_{k}\right\|_{2}^{\beta}\quad,\beta\in(0,2).$ (5)
For our implementation, we will consider $\beta=1$ and will write $S_{\mathrm{E}}^{(1)}(P,\mathbf{y})$ simply as $S_{\mathrm{E}}(P,\mathbf{y})$.
Note that, SRs such as energy scores can be expressed as $S(P_{\theta},\mathbf{y})=\mathbb{E}_{X,X^{\prime}\sim P_{\theta}}g(X,X^{\prime},\mathbf{y})$. Moreover, suppose the simulator model $P_{\theta}$ can also be represented as $X=h_{\theta}(Z)$, with $Z\sim Q$ where, the distribution $Q$ is independent of $\theta$. It has been shown in Pacchiardi et al. (2024b) that, if both $g$ and $h_{\theta}$ differentiable then an interchange of expectation and gradient step enables unbiased estimation of the gradient of the SR $S(P_{\theta},\mathbf{y})$ by using some random draws of $z_{i}\sim Q$ for $i=1,2,\ldots m$. This property is particularly useful in high-dimensional parameter spaces, where efficient exploration requires gradient information.
3.2 Prequential scoring rule posterior
As mentioned before, for a complex MDP we may not have the analytical form for the conditional probabilities $P_{\theta_{st}}$ and $P_{\theta_{r}}$ used in equation (1) to define the loglikelihood function. Hence combining the idea of scoring posterior explained in the previous section with the idea of prequential (predictive sequential) SR proposed in Pacchiardi et al. (2024a), here we propose prequential scoring rule posterior. For prequential SR, the original SRs are used to assess the goodness of fit of the one step conditional distributions using the observations at the next step given the observations at the previous step. Hence the prequential SR for a parametric MDP ($\mathbb{M}_{\theta}$) based on $H_{T}$ using a scoring rule $S$ would be given by,
$\operatorname{\mathcal{PS}}(\mathbb{M}_{\theta},H_{T})=\sum_{t=1}^{T}\left(S({P}_{\theta_{st}}(.|s_{t},a_{t}),s_{t+1})+S({P}_{\theta_{r}}(.|s_{t},a_{t}),r_{t+1})\right).$ (6)
When the scoring rule $S$ is (strictly) proper then the prequential scoring rule $\operatorname{\mathcal{PS}}(\mathbb{M}_{\theta},H_{T})$ is also (strictly) proper for the Markovian conditional distributions over the next state and reward given the previous state-action pair as shown in Theorem 2 of Pacchiardi et al. (2024a). If the true MDP belongs in class of MDPs =$\mathcal{M}=\{\mathbb{M}_{\theta}:\theta\in\Theta\}$ with the true parameter value being $\theta_{0}$, then the minimizer of the expected prequential scoring rule,
$\theta^{*}=\arg\min_{\theta}\mathbb{E}_{H_{T}\sim\mathbb{M}_{0}}\operatorname{\mathcal{PS}}(\mathbb{M}_{\theta},H_{T})$ (7)
would coincide with the true parameter $\theta_{0}$.
We notice, if we have access to the analytical form of $P_{\theta_{st}}$ and $P_{\theta_{r}}$, then using negative log-score the above equation reduces to equation (1).
If we can estimate the SR $S$ as shown in equation (5) unbiasedly, then the prequential SR $\operatorname{\mathcal{PS}}$ can also be estimated unbiasedly by generating $m$ simulations of next states and reward until time $T$, conditioned on previously observed state and action. Further in many MDPs, the reward distribution is chosen deterministically based on the problem and it is considered as a constant function of state and action. Hence, without loss of generality, in those cases, our $\operatorname{\mathcal{PS}}$ simplifies into
$\operatorname{\mathcal{PS}}(\mathbb{M}_{\theta},H_{T})=\sum_{t=1}^{T}S({P}_{\theta_{st}}(.|s_{t},a_{t}),s_{t+1}),$ (8)
and $\theta$ only contains parameters $\theta_{st}$. Based on this, we can define our prequential scoring rule posterior on the parameter $\theta$ of the parametric MDP as,
$\pi_{\operatorname{\mathcal{PS}}}\left(\theta\mid H_{T}\right)\propto\pi(\theta)\exp\left\{-w\operatorname{\mathcal{PS}}(\mathbb{M}_{\theta},H_{T})\right\},$ (9)
when $\pi(\theta)$ is the prior distribution on our parameters. Note that choosing values of $w$ other than $1$ can be interpreted as a form of annealing (heating it for $w<1$ and cooling it for $w>1$, we have taken $w=1$ for our experiments) applied to the target posterior.
Asymptotic properties of the prequential scoring rule posterior
Under the assumption of finite action space ($\mathcal{A}$ is countable and $|\mathcal{A}|=n_{a}<\infty$), we can also define prequential scoring rules for each of the actions as following,
$\operatorname{\mathcal{PS}}^{a}(\mathbb{M}_{\theta},H_{T})=\sum_{t=1}^{T}S({P}_{\theta_{st}}(.|s_{t},a_{t}),s_{t+1})\mathbb{I}_{a_{t}=a},$
where $\operatorname{\mathcal{PS}}(\mathbb{M}_{\theta},H_{T})=\sum_{a\in\mathcal{A}}\operatorname{\mathcal{PS}}^{a}(\mathbb{M}_{\theta},H_{T})$. Based on this, we can rewrite our prequential scoring rule posterior on the parameter $\theta$ of the parametric MDP as,
$\pi_{\operatorname{\mathcal{PS}}}\left(\theta\mid H_{T}\right)\propto\pi(\theta)\exp\left\{-w\sum_{a\in\mathcal{A}}\operatorname{\mathcal{PS}}^{a}(\mathbb{M}_{\theta},H_{T})\right\},$ (10)
when $\pi(\theta)$ is the prior distribution on our parameters. Further we denote, $T_{a}$ as the set consisting of timepoints when action $a$ was taken until time $T$, i.e., $T_{a}=\{t\leq T:a_{t}=a\}$ for each $a\in\mathcal{A}$. Assuming mild conditions on each of the action-specific Prequential SR loss functions $\operatorname{\mathcal{PS}}^{a}$ for $a\in\mathcal{A}$, the Prequential SR posterior satisfies a Bernstein-von Mises (BvM) theorem for finite action spaces. This theorem ensures the asymptotic normality of the corresponding posterior, with results analogous to the asymptotic behavior of the standard Bayesian posterior.
Next, we will prove the BvM theorem under some assumptions. Before stating the assumptions, we first clarify some mathematical notations. We assume $\Theta\subseteq\mathbb{R}^{d}$ and the minimizer of the prequential scoring rule $\theta^{*}\in E$, where $E\subseteq\Theta$ is open (in $\mathbb{R}^{d}$ ) and bounded. The prior $\pi:\Theta\rightarrow\mathbb{R}$ is a probability density function with respect to Lebesgue measure such that $\pi$ is continuous at $\theta^{*}$ and $\pi\left(\theta^{*}\right)>0$. For the sake of notational simplicity, from now on we will use $\operatorname{\mathcal{PS}}_{T}(\theta)$ to denote the prequential scoring rule $\operatorname{\mathcal{PS}}(\mathbb{M}_{\theta},H_{T})$ between the MDP $\mathbb{M}_{\theta}$ and $H_{T}$ and similarly $\operatorname{\mathcal{PS}}^{a}_{T}(\theta)$ to denote $\operatorname{\mathcal{PS}}^{a}(\mathbb{M}_{\theta},H_{T})$ for each action $a$. We denote the gradient, the matrix of second order derivatives and the tensor of third order derivatives of a function $f(\theta)$ with respect to $\theta$ as $f^{\prime}(\theta)=\left(\frac{\partial f(\theta)}{\partial\theta^{i}}(\theta)\right)_{i=1}^{d}\in\mathbb{R}^{d}$, $f^{\prime\prime}(\theta)=\left(\frac{\partial f}{\partial\theta^{i}\partial\theta^{j}}(\theta)\right)_{i,j=1}^{d}\in\mathbb{R}^{d^{2}}$ and $f^{\prime\prime\prime}(\theta)=\left(\frac{\partial f}{\partial\theta^{i}\partial\theta^{j}\partial\theta_{k}}(\theta)\right)_{i,j,k=1}^{d}\in\mathbb{R}^{d^{3}}$ respectively. For a given $\theta\in\mathbb{R}^{d}$ and $r>0$, we would denote the open ball of radius $r$ around $\theta$ as $B_{r}(\theta)=\{\theta^{\prime}\in\mathbb{R}^{d}:|\theta^{\prime}-\theta|<r\}$, where ||$\cdot$|| ⋅ | stands for Euclidean norm.
We first prove Lemma 1 under the following assumptions, which guarantee the existence of a sequence of consistent estimators of the true model parameter that converges to the expected SR minimizer defined in equation (7). It is important to note that Assumption A3 is a rather strong assumption regarding the convexity of each of the action specific prequential SRs. However, it is sufficient to verify Assumption A2 in conjunction with Assumption A1 to show the consistency.
  1. A1
    $\operatorname{\mathcal{PS}}_{T}^{a}:\Theta\rightarrow\mathbb{R}$ has continuous and uniformly bounded third derivatives on $E$, and if $T\rightarrow\infty$ then $\frac{\operatorname{\mathcal{PS}}_{T}^{a}}{|T_{a}|}\rightarrow\xi^{a}$ pointwise for some $\xi^{a}:\Theta\rightarrow\mathbb{R}$ with positive definite second derivative at $\theta^{*}$, denoted as $\operatorname{\Gamma}^{a}_{\star}$, $\forall a\in\mathcal{A}$.
  2. A2
    For some compact set $K\subseteq E$ with $\theta^{*}$ in its interior, $\xi^{a}(\theta)>\xi^{a}\left(\theta^{*}\right)$ for all $\theta\in K\backslash\left\{\theta^{*}\right\}$ and $\liminf_{T}\inf_{\theta\in\Theta\backslash K}\operatorname{\mathcal{PS}}_{T}^{a}(\theta)>\operatorname{\mathcal{PS}}_{T}^{a}\left(\theta^{*}\right)$, $\forall a\in\mathcal{A}$.
  3. A3
    $\operatorname{\mathcal{PS}}_{T}^{a}$ is convex and $\xi^{a\prime}\left(\theta^{*}\right)=0$, $\forall a\in\mathcal{A}$.
Lemma 1
If A1 and (A2 or A3) holds for all actions $a\in\mathcal{A}$, then the total prequential Scoring Rule loss, $\operatorname{\mathcal{PS}}_{T}(\theta)=\sum_{a\in\mathcal{A}}\operatorname{\mathcal{PS}}_{T}^{a}(\theta)$ also satisfies these conditions and there exists a sequence $\theta_{T}$ that converges to $\theta^{*}$ almost surely as $T\rightarrow\infty$.
We do not provide a proof as a similar result for generalized posteriors is established in Theorem 5 of Miller (2021). Note that, if the above conditions hold for each action-specific prequential SR $\operatorname{\mathcal{PS}}^{a}$, then the conditions are also satisfied by the summation $\operatorname{\mathcal{PS}}$. Consequently, Theorem 5 in Miller (2021) can be extended to prove the existence of a sequence of consistent estimators of $\theta^{*}$ as stated in Lemma 1. This result is analogous to the consistency of maximum likelihood estimators (MLE) and by setting $S$ as the log score, the original result can be recovered.
Next we prove the BvM type result for prequential scoring rules for MDPs under three additional assumptions. Note that assumption A4 ensures a polynomial approximation of the prequential SRs through a Taylor series expansion. Whereas assumption A6 ensures that all the actions are well explored in the training dataset. Under these assumptions, we formally state the theorem bellow.
  1. A4
    $\operatorname{\mathcal{PS}}_{T}^{a}$ can be represented as
    $\frac{1}{|T_{a}|}\operatorname{\mathcal{PS}}_{T}^{a}(\theta)=\frac{1}{|T_{a}|}\operatorname{\mathcal{PS}}_{T}^{a}\left(\theta_{T}\right)+\frac{1}{2}\left(\theta-\theta_{T}\right)^{\mathrm{T}}\operatorname{\Gamma}_{T}^{a}\left(\theta-\theta_{T}\right)+r_{T}^{a}\left(\theta-\theta_{T}\right),$
    where $\theta_{T}\in\mathbb{R}^{d}$ is such that $\theta_{T}\rightarrow\theta^{*}$ and $\operatorname{\Gamma}_{T}^{a}\in\mathbb{R}^{d\times d}$ is symmetric such that $\operatorname{\Gamma}_{T}^{a}\rightarrow\operatorname{\Gamma}_{*}^{a}$ as $|T_{a}|\rightarrow\infty$, $\forall a\in\mathcal{A}$. Further, there exist $\varepsilon_{0},c_{0}>0$ such that for all $T$ sufficiently large, for all $x\in B_{\varepsilon_{0}}(0)$, $\forall a\in\mathcal{A}$, $r_{T}^{a}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ satisfies $\left|r_{T}^{a}(x)\right|\leq c_{0}|x|^{3}$.
  2. A5
    For any $\varepsilon>0,\liminf_{n}\inf_{\theta\in B_{\varepsilon}\left(\theta_{T}\right)^{c}}\frac{1}{|T_{a}|}\left(\operatorname{\mathcal{PS}}_{T}^{a}(\theta)-\operatorname{\mathcal{PS}}_{T}^{a}\left(\theta_{T}\right)\right)>0$ $\forall a\in\mathcal{A}$.
  3. A6
    $\lim_{\begin{subarray}{c}|T_{a}|\rightarrow\infty\\ T\rightarrow\infty\end{subarray}}\frac{|T_{a}|}{T}=w_{a}\in(0,1)$, $\forall a\in\mathcal{A}$.
Theorem 2
[…]
that is, $\pi_{\operatorname{\mathcal{PS}}}(\theta|H_{T})$ concentrates at $\theta^{*}$,
$z_{T}\sim\frac{\exp\left(-\operatorname{\mathcal{PS}}_{T}\left(\theta_{T}\right)\right)\pi\left(\theta^{*}\right)}{\left|\operatorname{det}\operatorname{\Gamma}_{*}\right|^{1/2}}\left(\frac{2\pi}{T}\right)^{D/2}$ (12)
as $T\rightarrow\infty$ (Laplace approximation), with $\operatorname{\Gamma}_{*}=\sum_{a}w_{a}\operatorname{\Gamma}_{*}^{a}$; and letting $q_{T}$ be the density of $\sqrt{T}\left(\theta-\theta_{T}\right)$ when $\theta\sim\pi_{\operatorname{\mathcal{PS}}}\left(\theta\mid H_{T}\right)$,
$\int_{\mathbb{R}^{d}}\left|q_{T}(x)-\mathcal{N}\left(x\mid 0,\operatorname{\Gamma}_{*}^{-1}\right)\right|dx{\longrightarrow}0$ (13)
as $T\rightarrow\infty$, that is, $q_{T}$ converges almost surely to $\mathcal{N}\left(0,\operatorname{\Gamma}_{*}^{-1}\right)$ in total variation.
The above theorem implies that, as the number of observations, $T$, increases, the scoring rule posterior derived from these observations increasingly resembles a Gaussian distribution. This posterior distribution becomes progressively more concentrated around the expected scoring rule minimizer. We provide the full proof of Theorem 2 in Appendix A.1.
3.3 Sequential Monte Carlo with gradient based kernel
To sample from the prequential scoring rule posterior, here we propose to use a sequential Monte Carlo (Del Moral et al., 2006) scheme after every (or some) episode of interaction with the environment.
Sequential Monte Carlo (SMC)
Sequential Monte Carlo (SMC) refers to a class of algorithms which aim to represent a probability distribution through a set of weighted particles. Let $\{\pi_{k}\}$, for $k=1,2,\ldots$, to be the sequence of scoring rule posteriors using data from all the episodes up to the $k$-th episode. For each $k$, we use SMC to sample from the corresponding $\pi_{k}$, initializing the SMC with the posterior samples generated from the scoring rule posterior of the $(k-1)$-th episode. Hence the prior distribution used at the $k$-th episode to define the scoring rule posterior is $\pi_{k-1}$. To ensure a smooth transition from the prior ($\pi_{k-1}$) to the target distribution ($\pi_{k}$) at each episode, we introduce a series of intermediate target distributions. These distributions follow a geometric path, defined as ${\pi}^{(l)}_{k}(\theta)\propto\pi^{\alpha_{k,l}}_{k}(\theta)\pi_{k-1}^{1-\alpha_{k,l}}(\theta)$ with $10\leq\alpha_{k,1}<\ldots\alpha_{k,L}=$ 1 as proposed by Gelman and Meng (1998). The sequence of temperatures $\{\alpha_{k,l}\}_{l=1}^{L}$ at the $k$-th episode can be determined adaptively based on effective sample size (ESS)(Beskos et al., 2015) or conditioned effective sample size (CESS)(Zhou et al., 2016). Note that the prior distribution $\pi_{0}$ for the first episode, preferably an informative one, needs to be chosen.
The core of SMC relies on importance sampling across the particle trajectories. However, due to the high computational cost associated with sequential importance sampling, SMC introduces a series of backward kernels, used alongside forward kernels. Starting from an initial set of weighted particles drawn from a proposal distribution, each SMC iteration propagates these particles toward the target distribution via a forward kernel. The resulting trajectories are reweighted. The particles are then resampled according to these updated (normalized) weights. Note that we used an MCMC kernel with an invariant distribution matching the target posterior as the forward kernel and the corresponding time-reversal kernel as the backward kernel as outlined in Section 3.3.2.3 in Del Moral et al. (2006).
Before providing the details of the gradient based kernel, we first argue how we can easily derive an unbiased estimate of the prequential scoring rule. As discussed in Section 3.1, the prequential SR $S({P}_{\theta_{st}}(.|s_{t},a_{t}),s_{t+1})$ can also be expressed as an expectation over samples from $P_{\theta_{st}}$, conditioned on $(s_{t},a_{t})$ for all $t=1,2,\ldots T$. Whenever $S$ is differentiable with respect to $\theta$, an unbiased estimator of the gradient of the total prequential SR $\operatorname{\mathcal{PS}}_{T}(\theta)$ can be obtained using random samples from an auxiliary distribution such as a Gaussian or uniform distribution that is independent of $\theta$ using equation (5). Next, we describe the adSGRLD kernel used as the forward kernel for SMC.
Adjusted stochastic gradient Riemannian Langevin dynamic
Suppose we wish to sample from a distribution with pdf $p(\theta)=(1/Z)\exp(-U(\theta))$, $\theta\in\mathbb{R}^{d}$; where $Z$ is a normalizing constant and $U(\theta)$, the negative log-likelihood equivalent, is known as the potential energy. Standard SGLD is based on the Overdamped Langevin Diffusion, represented by the following stochastic differential equation which has a stationary distribution $p\left(\theta\right)$:
$d\theta(u)=-\frac{1}{2}\nabla_{\theta}U(\theta(u))du+dB_{u},$
where $B_{u}$ is a Brownian motion. To sample from the target distribution using the above SDE, we often use numerical approximation schemes like the Euler-Maruyama discretisation, which leads to the following update rule:
$\theta_{u+1}\leftarrow\theta_{u}-\frac{\epsilon}{2}\widehat{\nabla_{\theta}}U(\theta_{u})+\sqrt{\epsilon_{u}}W,$
where $W$ is a d-dimensional standard normal random vector; $\widehat{\nabla_{\theta}}U(\theta_{u})$ is an unbiased estimator of the gradient of $U(\theta_{u})$ and $\{\epsilon_{u}\}$ is a sequence of discretization step sizes satisfying the conditions, $\sum_{u=1}^{\infty}\epsilon_{u}=\infty$ and $\sum_{u=1}^{\infty}\epsilon_{u}^{2}<\infty$.
Although Welling and Teh (2011) have shown that SGLD yields samples from the target posterior when using a sequence $\epsilon_{u}$ converging to 00, in practice, $\epsilon_{u}$ seldom converges to 00, leading to bias due to Euler-Maruyama discretization. Hence, to ensure random sampling with minimal bias even when using a noisy estimate of the gradient, adaptive Langevin dynamics has been proposed in Jones and Leimkuhler (2011) which was later adapted for Bayesian inference (Ding et al., 2014) and likelihood-free inference (Pacchiardi et al., 2024b). The algorithm, referred to as adaptive stochastic gradient Langevin dynamics (adSGLD), runs on an augmented space $(\theta,p,\eta)$, where $\theta$ represents the parameter of interest, $p\in\mathbb{R}^{d}$ represents the momentum and $\eta$ represents an adaptive thermostat controlling the mean kinetic energy $\frac{1}{d}\mathbb{E}(p^{T}p)$.
Choosing an appropriate sequence of step sizes is crucial for effectively exploring the parameter space, especially in high dimensions. For example if different components of $\theta$ have values on different scales or the components are highly correlated, a poor choice of step sizes can result in slow mixing, negatively impacting the performance of the SMC sampler. To address this issue, several preconditioning schemes (Girolami and Calderhead, 2011) have been proposed in the literature. For instance, the Riemann manifold Metropolis-adjusted Langevin algorithm uses a positive definite matrix $G(\theta)$ to adaptively precondition the gradient. Then the SDEs are given by,
$d\bm{\theta}$ $=$ $G(\theta)\bm{p}du,$
$d\bm{p}$ $=$ $\left(-G(\bm{\theta})\nabla_{\bm{\theta}}U(\bm{\theta})-\bm{\eta}\bm{p}+\nabla_{\bm{\theta}}G(\bm{\theta})+G(\bm{\theta})(\bm{\eta}-G(\bm{\theta}))\nabla_{\bm{\theta}}G(\bm{\theta})\right)du+\sqrt{2}G(\bm{\theta})^{\frac{1}{2}}dB_{u},$
$d\bm{\eta}$ $=$ $\left(\frac{1}{d}\bm{p}^{T}\bm{p}-1\right)du$
where $G(\theta)$ is our preconditioning matrix, and $\nabla_{\bm{\theta}}G(\bm{\theta})$ is a vector with $i$-th element being $\sum_{j}\nabla_{\bm{\theta}_{j}}G_{ij}(\bm{\theta})$. $G(\bm{\theta})$ encodes geometric information of the potential energy $U(\bm{\theta})$, called Riemannian metric (Girolami and Calderhead, 2011), which are commonly defined by the Fisher information matrix.
If we assume $G(\bm{\theta})$ does not depend on $\bm{\theta}$, we notice that the third and fourth terms in the momentum update term vanishes and we end up with a SDE of the form
$d\bm{\theta}$ $=$ $G(\theta)\bm{p}du,$ (14)
$d\bm{p}$ $=$ $\left(-G(\bm{\theta})\nabla_{\bm{\theta}}U(\bm{\theta})-\bm{\eta}\bm{p}\right)du+\sqrt{2}G(\bm{\theta})^{\frac{1}{2}}dB_{u},$
$d\bm{\eta}$ $=$ $\left(\frac{1}{d}\bm{p}^{T}\bm{p}-1\right)du$
Using an Euler scheme with step size $\epsilon$, our algorithm becomes
Algorithm 2 Adjusted stochastic gradient Riemannian Langevin dynamic
Parameters $\epsilon,a$
Initialize $\bm{\theta}_{(0)}\in\mathbb{R}^{n},\mathbf{p}_{(0)}\sim\mathcal{N}(0,G(\bm{\theta}_{(0)}\epsilon\mathbf{I})$, and $\bm{\eta}_{(0)}=a$
for $t=1,2,\ldots$ do
     Evaluate $\nabla\tilde{U}(\bm{\theta}_{(t-1)})$
     $\mathbf{p}_{(t)}=\mathbf{p}_{(t-1)}-(\bm{\eta}_{(t-1)}\mathbf{p}_{(t-1)}+\nabla\tilde{U}(\bm{\theta}_{(t-1)})G(\bm{\theta}_{(t-1)}))\epsilon+\mathcal{N}(0,2a\epsilon G(\bm{\theta}_{(t-1)}))$
     $\bm{\theta}_{(t)}=\bm{\theta}_{(t-1)}+\mathbf{p}_{(t)}G(\bm{\theta}_{(t-1)})\epsilon$
     $\bm{\eta}_{(t)}=\bm{\eta}_{(t-1)}+(\frac{1}{d}\mathbf{p}_{(t)}^{T}\mathbf{p}_{(t)}-1)\epsilon$
end for
Note that, for our implementation, we have calculated the preconditioning matrix $G(\theta)$ using the gradient estimate as described in Chen et al. (2016).
4 ETS with generalized posterior
In Section 3, we have introduced a simulator model-based framework designed to obtain posterior samples when the model likelihood is unknown. The parallelizability of SMC samplers makes this algorithm scalable for higher dimensions and computationally efficient. Following this, we now present a result concerning the convergence of approximate policy iteration methods when integrated with ETS.
Evaluating a policy’s performance typically involves measuring regret, which requires calculating the value functions for both the optimal policy and the learned policy. This, in turn, depends on knowing the exact state transition probabilities and reward distributions of the true environment to compute the expectation. However, for many complex tasks, these exact distributions defining the environment’s dynamics are intractable, making it necessary to approximate the value functions from observed interaction data.
Therefore, we provide a convergence result for the ETS-based policy in terms of the difference between the action-value functions of the optimal policy and the learned policy. Additionally, we prove this result for a well-specified case where the expected scoring rule minimizer aligns with the true model parameters.
Theorem 3
Let $\mu_{1},\mu_{2},\ldots$ be the sequence of policies generated through ETS with an approximate policy iteration algorithm after $k$ episodes of interaction with the environment. Also, let $Q^{\bm{\theta^{(k)}}}_{\mu_{j}}$ denote the estimated $Q$ function for the policy $\mu_{j}$ with $\mu_{j+1}(s)=\arg\max_{a}Q^{\bm{\theta^{(k)}}}_{\mu_{j}}(s,a)$ for all $j=1,2,\ldots$ Assuming $Q_{\mu}^{\theta}$ to be Lipschitz continuous for all possible policies $\mu$ and $\theta\in\Theta$, for the case of a well-specified model we have
$||Q^{*}-Q^{\bm{\theta^{(k)}}}_{\mu_{j+1}}||_{\infty}\leq\gamma^{j}||Q^{*}-Q^{\bm{\theta^{(k)}}}_{\mu_{1}}||_{\infty}+\sum_{l=1}^{j}\gamma^{j-l+1}\zeta_{l}(k,n),$
where $\gamma\in[0,1]$ is a discounting factor used to define the value functions.
The above theorem suggests, irrespective of the number of episodes observed, the sequence of policies obtained from a policy iteration method integrated with ETS converges progressively toward optimal behavior as the iterations increase. The second term $\zeta_{l}(k,n)$ involves the samples drawn from the posterior of the model parameters obtained after observing up to $k$ episodes of interaction with the environment. As the episode count $k$ and the number of posterior samples $n$ increase, this term shrinks to 00 for any $l=1,2,\ldots$ according to Theorem 2 as the posterior distribution of the model parameters concentrates around the Expected SR minimizer. The detailed proof of the above theorem can be found in Appendix A.2.
5 Simulation studies
In this section, we demonstrate the application of several model-free policy learning algorithms integrated with ETS, comparing the performance of the ETS-integrated approach with that of the classical model-free method. We begin by presenting results for a finite action space problem, showing both well-specified and misspecified model cases. We then extend the analysis to a problem with continuous action space, where we focus on the misspecified model scenario.
5.1 Finite action MDP
Well-specified models
To demonstrate the Scoring rule Thompson Sampling using SMC, we use the ‘inverted pendulum’ task from Dimitrakakis and Tziortziotis (2013). The agent targets to keep the pendulum upright as long as possible by switching actions. State of the environment is defined as (angle, angular velocity) of the pendulum. There are three possible actions from each state. The actions are force (in Newtons) in a certain direction, the action space is $(+50,0,-50)$. The problem has 6 parameters: the pendulum mass, the cart mass, the pendulum length, the gravity, the amount of uniform noise, and the simulation time interval. In this environment, agent gets $+1$ reward for every balancing step. The actual environment parameters are $\theta_{0}=(2.0,8.0,0.5,9.8,10,0.01)$ and the simulator model ($P_{\theta}$) for the dynamics of the environment is known to us.
In our comparison, we evaluated the online performance of both the model-free and Bayesian model-based approaches in the pendulum domain. In the model-free approach, we updated our policy after each episode of interaction with the environment. The policy updates ceased once we reached the maximum reward, which is known to be $1000$ in our setup.
Figure 4: Both the policies are trained using the same dataset generated through a random policy. The learned policy is used to interact with the environment for a maximum of 1000 steps or until the pendulum falls for the first time. The same experiment is run independently 10 times and the average discounted return is plotted.
For policy learning, we adopted the least squares policy iteration (LSPI) method proposed by Lagoudakis and Parr (2003). LSPI is a model-free approximate policy iteration method that leverages the least squares temporal-difference learning algorithm to approximate the $Q$ function. Specifically, we approximated the $Q$ function using a linear combination of a $4\times 4$ grid of Gaussian radial basis functions in LSPI, with a learning rate of $\gamma=0.99$.
Additionally, for the Bayesian approach, we assumed a uniform prior over the interval $[0.5\theta_{0},5.0\theta_{0}]$ for the model parameters. We obtained samples from the posterior distribution after each episode using SMC with the adSGLD kernel (Jones and Leimkuhler, 2011) (refer to Table 1 for details of the hyperparameters). To define the intermediate target distributions between consecutive episodes, we chose the sequence of temperatures $\{\alpha_{k,l}\}_{l=1}^{L}$ at the $k$-th episode, such that the effective sample size (ESS) declines uniformly through out the SMC iterations. We used bisection method to find the temperature, $\alpha_{k,l}$ such that $\text{ESS}_{l}=\kappa\times\text{ESS}_{l-1}$, the parameter $\kappa$ for each experiment is mentioned in Appendix A.3. Once the ESS drops below half the original sample size, the particles are resampled. Here we used the zeroth order gradient (discussed in A.4) to compute an estimate of the gradient of the prequential loss. We used LSPI integrated with ETS to learn an optimal policy, as described in Section 2.1 where we updated the policy for a maximum of $J=30$ iterations in-between each episode.
Figure 4 illustrates that both the model-free and Bayesian model-based methods exhibit improved policy learning with increased training data. However, the Bayesian approach achieves optimal rewards much sooner compared to model-free training. This expedited convergence in Bayesian methods can be attributed to the fact that once the posterior distribution of the model parameters converges, simulated interactions closely resemble true interactions. With this concentrated posterior distribution, longer trajectories of simulated data can be generated, facilitating more accurate estimation of the value function. Consequently, when using these longer chains of simulated data to estimate the value function and perform LSPI to find an optimal policy, convergence occurs more rapidly due to the utilization of a more stable estimate of the value function.
Misspecified models
To demonstrate the application of ETS for the case where a simulator model is not known, we revisit the ‘inverted pendulum’ experiment with the previously mentioned parameters. Conditional GANs (Charlesworth and Montana, 2020; Zhao et al., 2021) have been widely adopted in the literature to model the dynamics. We have used a generative neural network, skipping the adversarial training, to model the difference between the next state and the current state conditioned on the current state and action as suggested by Nagabandi et al. (2018) and Deisenroth et al. (2013). Further, for the states $s$ that represent the angle, we have considered feeding in $(\sin(s),\cos(s))$ as inputs to the model. The model can be written as,
$f_{\theta}(s_{t},a_{t},z)=s_{t+1}-s_{t}$
where $z$ is a Gaussian noise and $f_{\theta}$ is defined by $3$ fully connected layers, $10$ neurons per layer and ‘swish’ activation functions (Ramachandran et al., 2017) following the architecture suggested by Chua et al. (2018).
For parameter inference we have used Sequential Monte Carlo (SMC) (implementation details can be found in Section A.3 in the Appendix). In high-dimensional parameter spaces, selecting a well-chosen initial sample from an informative prior is crucial for effective Bayesian inference. To do so, we ran the Adam optimizer (Kingma, 2014) with a learning rate of $0.001$ for $1000$ steps, saving the final $100$ points. The covariance of these optimized points was computed and used to add Gaussian noise, generating a total of $300$ particles to initialize the SMC. To define the intermediate target distributions between consecutive episodes, we chose the sequence of temperatures $\{\alpha_{k,l}\}_{l=1}^{L}$ at the $k$-th episode, such that the conditional effective sample size (CESS) (Zhou et al., 2016) stays constant through out the SMC iterations. We used bisection method to find the temperature, $\alpha_{k,l}$ such that $\text{ESS}_{l}=\kappa\times N$, where $N$ is the sample size and the parameter $\kappa$ for each experiment is mentioned in Appendix A.3. Once the ESS drops below $N/2$, the particles are resampled.
Figure 5: Both the policies are trained using the same dataset generated through a random policy. The learned policy is used to interact with the environment for a maximum of 1000 steps or until the pendulum falls for the first time. The same experiment is run independently 10 times and the average discounted return is plotted.
Figure 5 demonstrates that LSPI combined with ETS learns the optimal policy significantly faster than its model-free counterpart. Although the true dynamics of the environment were unknown, the generative neural network effectively learned the dynamics, accelerating the policy learning process. Note that the Theorem 2 on the consistency of the generalized posterior distribution and the Theorem 3 on the bound on the approximation error of the $Q$ function under ETS assume a unimodal posterior for model parameters. Whilst this assumption may not hold when using a generative neural network as a simulator model, we still demonstrate that ETS is more sample-efficient for simpler models.
SMC performance generally benefits from an increased number of particles, which raises computational costs. Hence, we used only $300$ particles for SMC and resampled $100$ particles out of them based on their weights for ETS. This streamlined approach still yielded promising results.
5.2 Continuous action MDP
When the action space is continuous, calculating the $Q$-function for all possible state-action pairs is impossible. A deterministic policy that maximizes the $Q$-function over the entire action space cannot be easily found. In such cases, a parameterized probabilistic policy is used, where $\mu_{\alpha}(a|s)$ denotes the probability of taking action $a$ when the environment is in state $s$, and $\alpha\in\mathbb{R}^{d^{\prime}}$ are the policy parameters. Usually the policy distribution is considered as a Gaussian distribution and $\alpha$ would contain the mean and covariance of the distribution. The agent here tries to update the policy parameters according based on the rewards collected during realtime interaction with the environment.
In classical policy gradient methods, a performance measure $J(\alpha)$ is computed from real interaction data, and the policy parameters are updated using gradient ascent according to:
$\alpha_{t+1}\leftarrow\alpha_{t}+\widehat{\nabla}J(\alpha_{t}),$
where ^$\widehat{\nabla}J(\alpha^{\prime})$ is an estimate of the gradient of $J$ with respect to $\alpha$ evaluated at $\alpha^{\prime}$.
In ETS, the performance measure of a policy is estimated from simulated interactions based on each posterior sample of the model parameters. Let $J^{\theta}(\alpha)$ represent the performance measure of the policy $\mu_{\alpha}$ based on interactions simulated from the MDP $\mathbb{M}_{\theta}$. The pooled estimate of $J$ for policy $\mu{\alpha}$, based on posterior samples $\bm{\theta}^{(k)}$, is given by:
[…]
Then, the policy parameters $\alpha$ are updated as
$\alpha_{t+1}\leftarrow\alpha_{t}+\widehat{\nabla}J^{\bm{\theta}^{(k)}}(\alpha_{t}).$ (16)
We perform the policy update based on the simulated trajectories until convergence and use the latest policy to interact with the environment in the next episode. Although we do not present theoretical results for ETS applied to continuous action spaces, our empirical findings align closely with the theoretical results we previously established for discrete action spaces.
Example
To demonstrate the ETS strategy on a problem with continuous action space for the case of a misspecified model, we choose the ‘Hopper’ experiment from the OpenAI Gymnasium package (Kwiatkowski et al., 2024). This problem involves moving a two-dimensional, single-legged structure forward by applying force to three joints connecting its four body parts. Actions here consist of torques (ranging from -1 Nm to 1 Nm) applied to each of the three joints, while the environment’s state is defined as a $12$-dimensional real-valued vector of joint angles, angular velocities, positions, and velocities of the body parts. The reward function incentivizes forward movement of the hopper while penalizing the application of excessive torque, which could destabilize the system.
We have used a similar neural network architecture as described in Sction 5.1, for our simulator model, using three fully connected layers with $20$ nodes each. For sampling from this posterior of the model parameters, we used SMC, as before, but to maintain computational efficiency we used only $45$ particles (this figure was chosen based on the runs with 100 and 500 particles, which were not providing any significant improvements while already being expensive) for this problem. As before, we ran the Adam optimizer (Kingma, 2014) with a learning rate of $0.001$ for $1000$ steps, saving the final $15$ points. The covariance of these optimized points was computed and used to add Gaussian noise, generating a total of $45$ particles to initialize the SMC. Similar to misspecified ‘inverted pendulum’ problem, we defined intermediate target distributions for SMC based on the CESS.
For policy learning, we applied the ‘REINFORCE’ (Williams, 1992) a foundational policy gradient method suitable for continuous action spaces, integrating ETS via equation (15). The policy parameters were then updated based on equation (16). Full implementation details of the experiment can be found in Section A.3 and A.5 in the Appendix. Given that with an increasing amount of data, the posterior distribution can become overly concentrated, we limited posterior updates and sampling to the first $15$ episodes. At episode $15$, we halted model training and proceeded with the standard REINFORCE updates to refine the policy during future interactions with the environment.
Figure 6: For visual clarity we have considered the moving average (lag 10) of episodic returns collected using model-free REINFORCE and ETS-integrated REINFORCE. The average of the returns over five different seeds are plotted for both cases.
In Fig. 6, we compare the performance of REINFORCE integrated with ETS versus the classical model-free approach, where we notice that REINFORCE combined with ETS quickly learns high-rewarding policies, outperforming its model-free counterpart in terms of sample-efficiency. Although model misspecification is present, the universal approximation theorem (Hornik et al., 1989; Hornik, 1991) suggests that this error can be minimized by carefully selecting the complexity of the neural networks. Hence the advantage of ETS-integrated policy learning arises from its generative neural network-based model-learning component, which accelerates policy convergence.
6 Conclusion
In this work, we have introduced a robust Bayesian framework for model-based reinforcement learning. A fully Bayesian treatment of model parameters traditionally requires a tractable likelihood function. This is often unavailable when the environment’s dynamics is learned through generative neural networks. To address this issue, we formulated a generalized posterior for the model parameters using prequential scoring rules based on a Markovian assumption on observed trajectories, enabling generalized Bayesian inference in the absence of a known likelihood. Additionally, we established a Bernstein-von Mises (BvM) type consistency result for the scoring rule-based generalized posterior in the discrete action setting, leaving the proof for continuous action space as future work.
For efficient sampling from this generalized posterior, we used SMC samplers with an adSGLD kernel to leverage gradient information for better exploration in high-dimensional parameter spaces. We further improve the kernel with gradient-based preconditioning matrices similar to those in the Adam optimizer to provide better guidance for sampling in high-dimensions. For policy learning, we extended classical Thompson sampling by incorporating full posterior samples for enhanced policy search, introducing the expected Thompson sampling (ETS) approach. Using the BvM result, we derived an error bound for approximating the action-value function when a policy iteration method is integrated with ETS in well-specified model settings.
To empirically evaluate the proposed method, we first compared classical model-free least squares policy iteration with the ETS-integrated Bayesian version on a simple inverted pendulum balancing task. In both well-specified and misspecified model settings, the ETS-integrated approach learned the optimal policy more quickly than the model-free baseline. We then applied ETS to a more complex problem of teaching a single-legged hopper to move forward without falling. Even with continuous action space and a misspecified model, the ETS-based approach discovered a high-reward policy significantly faster than the model-free approach.
In conclusion, we present a robust framework for policy learning that enables rapid identification of high-reward policies in complex tasks. This approach is especially relevant for applications in robotics and autonomous systems where agents must make quick decisions without the luxury of prolonged learning periods. However, we observed that using generative deep neural networks as a simulator model leads to a multimodal posterior for model parameters. This makes sampling difficult, especially in high dimensions. We are keen to explore if a shrinkage prior on the parameter space can address this issue. Besides, future directions of this work include investigating copula-based nonparametric methods for model inference, potentially mitigating the effects of model misspecification. Another promising avenue would be integrating Bayesian model inference and policy training under a unified objective, optimizing both processes concurrently.
Appendix A
A.1 Proof of BvM, Theorem 2
Here we extend the proof of Bernstein–von Mises (BvM) theorem for generalized posterior (Miller, 2021) for our setup.
Proof  Let us define a function $g_{T}(x):\Theta\rightarrow\mathbb{R}$ as following,
$g_{T}(x)$ $=\exp\left(-\left[\operatorname{\mathcal{PS}}_{T}\left(\theta_{T}+x/\sqrt{T}\right)-\operatorname{\mathcal{PS}}_{T}\left(\theta_{T}\right)\right]\right)\pi\left(\theta_{T}+x/\sqrt{T}\right)$
$=q_{T}(x)z_{T}e^{\operatorname{\mathcal{PS}}_{T}\left(\theta_{T}\right)}T^{D/2}\quad\text{where, $q_{T}(x)=\pi_{\operatorname{\mathcal{PS}}}\left(\theta_{T}+x/\sqrt{T}|H_{T}\right)T^{-D/2}$ };$ (17)
recalling that $z_{T}<\infty$ by assumption, and define
$g_{0}(x)=\exp\left(-\frac{1}{2}x^{\prime}\operatorname{\Gamma}_{*}x\right)\pi\left(\theta^{*}\right)\text{with $\operatorname{\Gamma}_{*}=\sum_{a}w_{a}\operatorname{\Gamma}_{*}^{a}$}$
If $\lambda^{a}$ is the smallest eigenvalue of $H_{*}^{a}$ for each action $a\in\mathcal{A}$, then let us consider $\alpha^{a}\in(0,\lambda^{a})$ and a $\epsilon^{*}>0$ small enough such that $\epsilon^{*}<\alpha^{a}/\left(2c_{0}\right)\ \forall a\in\mathcal{A},\epsilon^{*}<\epsilon_{0}$, and $\pi(\theta)\leq 2\pi\left(\theta^{*}\right)$ for all $\theta\in B_{2\epsilon^{*}}\left(\theta^{*}\right)$ (as $\pi$ is continuous at $\theta^{*}$ ). As $\liminf_{T}\inf_{\theta\in B_{\epsilon^{*}}\left(\theta_{T}\right)^{c}}\frac{1}{|T_{a}|}(\operatorname{\mathcal{PS}}_{T}^{a}(\theta)-\operatorname{\mathcal{PS}}_{T}^{a}\left(\theta_{T}\right))>0$ for any $\epsilon>0$ from assumption A5, we define $\delta_{a}=\liminf_{T}\inf_{\theta\in B_{\epsilon^{*}}\left(\theta_{T}\right)^{c}}\frac{1}{|T_{a}|}(\operatorname{\mathcal{PS}}_{T}^{a}(\theta)-\operatorname{\mathcal{PS}}_{T}^{a}\left(\theta_{T}\right))>0,\forall a\in\mathcal{A}$ and $\delta=\sum_{a\in\mathcal{A}}\frac{|T_{a}|}{T}\delta_{a}.$
Next we define two functions, $h_{T}(x):\mathbb{R}^{d}\rightarrow\mathbb{R}$ and $h_{0}(x):\mathbb{R}^{d}\rightarrow\mathbb{R}$, as following
$h_{T}(x)$ =$=\begin{cases}\exp\left(-\frac{1}{2}x^{\prime}A_{T}x\right)2\pi\left(\theta^{*}\right)&\text{ if }|x|<\varepsilon^{*}\sqrt{T},\\ e^{-T\delta/2}\pi\left(\theta_{T}+x/\sqrt{T}\right)&\text{ if }|x|\geq\epsilon^{*}\sqrt{T}\end{cases}$
$h_{0}(x)$ $=\exp\left(-\frac{1}{2}x^{\prime}A_{0}x\right)2\pi\left(\theta^{*}\right)$
where $A_{T}^{a}=\operatorname{\Gamma}_{T}^{a}-\alpha^{a}I_{d}$, $A_{0}^{a}=\operatorname{\Gamma}_{*}^{a}-\alpha^{a}I_{d}$, $A_{T}=\sum_{a}\frac{|T_{a}|}{T}A_{T}^{a}$ and $A_{0}=\sum_{a}w_{a}A_{0}^{a}$. We note that, if
  1. (a)
    $g_{T}\rightarrow g_{0}$ and $h_{T}\rightarrow h_{0}$ pointwise,
  2. (b)
    $\int h_{T}\rightarrow\int h_{0}$,
  3. (c)
    $g_{T}=\left|g_{T}\right|\leq h_{T}$ for all $T$ sufficiently large, and
  4. (d)
    $g_{T},g_{0},h_{T},h_{0}\in L^{1}$ for all $T$ sufficiently large.
then by the generalized dominated convergence theorem, $\int g_{T}\rightarrow\int g_{0}$ and $\int\left|g_{T}-g_{0}\right|\rightarrow 0$ (e.g., Folland, 2013, exercises 2.20, 2.21). Since $\int q_{T}=1$, by equation (A.1) we have
$e^{\operatorname{\mathcal{PS}}_{T}\left(\theta_{T}\right)}T^{d/2}z_{T}=\int g_{T}\longrightarrow\int g_{0}=\pi\left(\theta^{*}\right)\frac{(2\pi)^{d/2}}{\left|\operatorname{\Gamma}_{*}\right|^{1/2}},$ (18)
where $\left|\operatorname{\Gamma}_{*}\right|=\left|\operatorname{det}\operatorname{\Gamma}_{*}\right|$, and hence,
$z_{T}\sim\frac{e^{-\operatorname{\mathcal{PS}}_{T}\left(\theta_{T}\right)}\pi\left(\theta^{*}\right)}{\left|\operatorname{\Gamma}_{*}\right|^{1/2}}\left(\frac{2\pi}{T}\right)^{d/2}\quad\text{as $T\rightarrow\infty$}.$
This proves equation (12) in our Theorem 2.
As $T\rightarrow\infty$, for any sequence $b_{T}\rightarrow b_{0}\in\mathbb{R}$, we have $\int\left|b_{T}g_{T}-b_{0}g_{0}\right|\rightarrow 0$, since (as $\int\left|g_{T}-g_{0}\right|\rightarrow 0$ and using (d))
$\int\left|b_{T}g_{T}-b_{0}g_{0}\right|\leq\int\left|b_{T}g_{T}-b_{T}g_{0}\right|+\int\left|b_{T}g_{0}-b_{0}g_{0}\right|=\left|b_{T}\right|\int\left|g_{T}-g_{0}\right|+\left|b_{T}-b_{0}\right|\int\left|g_{0}\right|\longrightarrow 0.$
Letting $1/b_{T}=e^{\operatorname{\mathcal{PS}}_{T}\left(\theta_{T}\right)}T^{d/2}z_{T}$ and $1/b=\pi\left(\theta^{*}\right)\frac{(2\pi)^{d/2}}{\left|\operatorname{\Gamma}_{*}\right|^{1/2}}$, we have $b_{T}\rightarrow b$ by equation (18), and
$\int_{\mathbb{R}^{d}}\left|q_{T}(x)-\mathcal{N}\left(x\mid 0,\operatorname{\Gamma}_{*}^{-1}\right)\right|dx\underset{T\rightarrow\infty}{\longrightarrow}0,$
which proves equation (13) in our Theorem 2 and the concentration of $\pi_{T}$ around $\theta^{*}$ in equation (11) follows from the Lemma 28 in Miller (2021). To conclude the proof of the Theorem, next we show that the conditions mentioned above (a)–(d) hold.
Proof of (a)
Let us fix some $x\in\mathbb{R}^{d}$. For sufficiently large $T$, we have $|x|<\epsilon^{*}\sqrt{T}$, then,
$h_{T}(x)=\exp\left(-\frac{1}{2}x^{\prime}(\sum_{a\in\mathcal{A}}\frac{|T_{a}|}{T}A^{a}_{T})x\right)2\pi(\theta^{*})\longrightarrow\exp\left(-\frac{1}{2}x^{\prime}(\sum_{a\in\mathcal{A}}w_{a}A^{a}_{0})x\right)2\pi(\theta^{*})$
since $A^{a}_{T}\rightarrow A^{a}_{0}$ and $\frac{|T_{a}|}{T}\rightarrow w_{a}$ as $T\rightarrow\infty$ by assumptions A4 and A6 respectively and $|\mathcal{A}|<\infty$
Now, for $g_{T}$, first note that $\pi\left(\theta_{T}+x/\sqrt{T}\right)\rightarrow\pi\left(\theta_{0}\right)$ since $\pi$ is continuous at $\theta_{0}$ and $\theta_{T}\rightarrow\theta_{0},x/\sqrt{T}\rightarrow 0$. By assumption A4 for all action $a\in\mathcal{A}$, we have
$\operatorname{\mathcal{PS}}^{a}_{T}(\theta_{T}+x/\sqrt{T})-\operatorname{\mathcal{PS}}^{a}_{T}(\theta_{T})=\frac{1}{2}\frac{|T_{a}|}{T}x^{\prime}\operatorname{\Gamma}^{a}_{T}x+|T_{a}|r^{a}_{T}(x/\sqrt{T})\longrightarrow\frac{1}{2}w_{a}x^{\prime}\operatorname{\Gamma}^{a}_{*}x,$ (19)
since $\operatorname{\Gamma}^{a}_{T}\rightarrow\operatorname{\Gamma}^{a}_{*}$ and for all $T$ sufficiently large (to ensure that $|x/\sqrt{T}|<\varepsilon_{0}$ and the assumed bound on $r^{a}_{T}$ holds),
$\left||T_{a}|r^{a}_{T}(x/\sqrt{T})\right|\leq|T_{a}|c_{0}|x/\sqrt{T}|^{3}=\frac{|T_{a}|}{T}\frac{c_{0}|x|^{3}}{\sqrt{T}}\rightarrow 0$
as $T\rightarrow\infty$. Next, let us consider
$g_{T}(x)$ $=\exp[-\sum_{a}(\operatorname{\mathcal{PS}}^{a}_{T}(\theta_{T}+x/\sqrt{T})-\operatorname{\mathcal{PS}}^{a}_{T}(\theta_{T}))]\pi(\theta_{T}+x/\sqrt{T})$
$=\prod_{a}[\exp[-(\operatorname{\mathcal{PS}}^{a}_{T}(\theta_{T}+x/\sqrt{T})-\operatorname{\mathcal{PS}}^{a}_{T}(\theta_{T}))](\pi(\theta_{T}+x/\sqrt{T}))^{\frac{1}{|\mathcal{A}|}}]$
[…]
$=\exp\left(-\frac{1}{2}x^{\mathrm{T}}\operatorname{\Gamma}_{*}x\right)\pi\left(\theta^{*}\right)=g_{0}(x).$
Hence, $g_{T}(x)\rightarrow g_{0}(x)$ as $T\rightarrow\infty$, which completes the proof of (a).
Proof of (b)
By the definition of $h_{T}$ and $A_{T}=\sum_{a}\frac{|T_{a}|}{T}A_{T}^{a}$, letting $B_{T}=B_{\varepsilon^{*}\sqrt{T}}(0)$,
[…]
Since $A^{a}_{T}\rightarrow A^{a}_{0}$ as $T\rightarrow\infty$ and $A^{a}_{0}$ is positive definite, then for all $T$ sufficiently large, $A^{a}_{T}$ is also positive definite for all $a\in\mathcal{A}$. Then the first term equals
$2\pi\left(\theta_{0}\right)\frac{(2\pi)^{d/2}}{\left|\sum_{a}\frac{|T_{a}|}{T}A^{a}_{T}\right|^{1/2}}{P}\left(\left|\left(\sum_{a}\frac{|T_{a}|}{T}A^{a}_{T}\right)^{-1/2}Z\right|<\varepsilon^{*}\sqrt{T}\right)\longrightarrow 2\pi\left(\theta_{0}\right)\frac{(2\pi)^{d/2}}{\left|\sum_{a}w_{a}A^{a}_{0}\right|^{1/2}}=\int h_{0},$
where $Z\sim\mathcal{N}(0,I)$.
The second term goes to zero, since it is nonnegative and upper bounded by
[…]
using the fact that $\pi\left(\theta_{T}+x/\sqrt{T}\right)T^{-d/2}$ is the density of $X=\sqrt{T}\left(\theta-\theta_{T}\right)$ when $\theta\sim\pi$.
Proof of (c)
We consider $T$ sufficiently large such that $\left|\theta_{T}-\theta^{*}\right|<\varepsilon^{*}$ (using Lemma  1), the bound on $r^{a}_{T}$ applies, and $\inf_{\theta\in B_{\varepsilon^{*}}\left(\theta_{T}\right)^{c}}\frac{1}{|T_{a}|}(\operatorname{\mathcal{PS}}^{a}_{T}(\theta)-\operatorname{\mathcal{PS}}^{a}_{T}\left(\theta_{T}\right))>\delta_{a}/2$ for all $a\in\mathcal{A}$ by assumption A5.
If $|x|\geq\varepsilon^{*}\sqrt{T}$, then $\left(\operatorname{\mathcal{PS}}^{a}_{T}\left(\theta_{T}+x/\sqrt{T}\right)-\operatorname{\mathcal{PS}}^{a}_{T}\left(\theta_{T}\right)\right)>|T_{a}|\delta_{a}/2$ for all $a\in\mathcal{A}$, and thus,
$\operatorname{\mathcal{PS}}_{T}\left(\theta_{T}+x/\sqrt{T}\right)-\operatorname{\mathcal{PS}}_{T}\left(\theta_{T}\right)>\frac{T}{2}\sum_{a}\frac{|T_{a}|}{T}\delta_{a}=\frac{T\delta}{2},$
which implies,
$g_{T}(x)\leq e^{-T\delta/2}\pi\left(\theta_{T}+x/\sqrt{T}\right)=h_{T}(x).$
On the other hand, if $|x|<\varepsilon^{*}\sqrt{T}$, then $\pi\left(\theta_{T}+x/\sqrt{T}\right)\leq 2\pi\left(\theta_{0}\right)$ (by our choice of $\varepsilon^{*}$, since $|(\theta_{T}+x/\sqrt{T})-\theta_{0}|\leq|\theta_{T}-\theta_{0}|+|x/\sqrt{T}|<2\varepsilon^{*})$. Further, $\left||T_{a}|r^{a}_{T}(x/\sqrt{T})\right|\leq\frac{|T_{a}|}{T}\frac{c_{0}|x|^{3}}{\sqrt{T}}\leq\frac{|T_{a}|}{T}c_{0}\varepsilon^{*}|x|^{2}\leq\frac{1}{2}\frac{|T_{a}|}{T}\alpha^{a}|x|^{2}$, by the fact that $|x/\sqrt{T}|<\varepsilon^{*}<\varepsilon_{0}$ and $\varepsilon^{*}<\alpha^{a}/\left(2c_{0}\right)$ for all $a\in\mathcal{A}$. Using the assumption A4 we get,
$\operatorname{\mathcal{PS}}_{T}\left(\theta_{T}+x/\sqrt{T}\right)-\operatorname{\mathcal{PS}}_{T}\left(\theta_{T}\right)$ $=\sum_{a}\left[\frac{1}{2}\frac{|T_{a}|}{T}x^{\prime}\operatorname{\Gamma}^{a}_{T}x+|T_{a}|r^{a}_{T}(x/\sqrt{T})\right]$
$\geq\sum_{a}\left[\frac{1}{2}\frac{|T_{a}|}{T}x^{\prime}\operatorname{\Gamma}^{a}_{T}x-\frac{1}{2}\frac{|T_{a}|}{T}\alpha^{a}x^{\prime}x\right]$
$=\frac{1}{2}x^{\prime}\left(\sum_{a}\frac{|T_{a}|}{T}(\operatorname{\Gamma}^{a}_{T}-\alpha^{a}I_{d})\right)x=\frac{1}{2}x^{\prime}\left(\sum_{a}\frac{|T_{a}|}{T}A^{a}_{T}\right)x=\frac{1}{2}x^{\prime}A_{T}x,$
implying,
$g_{T}(x)\leq\exp\left(-\frac{1}{2}x^{\prime}A_{T}x\right)2\pi\left(\theta_{0}\right)=h_{T}(x).$
Proof of (d)
Since $\operatorname{\Gamma}^{a}_{*}$ and $A^{a}_{0}$ are positive definite for all $a\in\mathcal{A}$, $\int g_{0}$ and $\int h_{0}$ are finite. By (b) and (c), since $\int h_{T}\rightarrow\int h_{0}<\infty$, we have $\int g_{T}\leq\int h_{T}<\infty$ for all $T$ sufficiently large. Measurability of $g_{T}$ and $h_{T}$ follows from measurability of $\operatorname{\mathcal{PS}}_{T}$ and $\pi$. This concludes our proof.
A.2 Approximation error of the Q function
Proof  For a fixed episode $k$, we have
$||Q^{*}-Q^{\bm{\theta^{(k)}}}_{\mu_{j+1}}||_{\infty}$ $=||Q^{*}-Q^{\theta^{*}}_{\mu^{*}_{j+1}}+Q^{\theta^{*}}_{\mu^{*}_{j+1}}-Q^{\bm{\theta^{(k)}}}_{\mu_{j+1}}||_{\infty}$
$\leq||Q^{*}-Q^{\theta^{*}}_{\mu^{*}_{j+1}}||_{\infty}+||Q^{\theta^{*}}_{\mu^{*}_{j+1}}-Q^{\theta^{*}}_{\mu_{j+1}}+Q^{\theta^{*}}_{\mu_{j+1}}-\frac{1}{n}\sum_{i=1}^{n}Q^{{\theta}_{ki}}_{\mu_{j+1}}||_{\infty}$
$\leq\gamma||Q^{*}-Q^{\theta^{*}}_{\mu^{*}_{j}}||_{\infty}+||Q^{\theta^{*}}_{\mu^{*}_{j+1}}-Q^{\theta^{*}}_{\mu_{j+1}}||_{\infty}+\frac{1}{n}\sum_{i=1}^{n}||Q^{\theta^{*}}_{\mu_{j+1}}-Q^{{\theta}_{ki}}_{\mu_{j+1}}||_{\infty}$
[…]
$=\gamma||Q^{*}-Q^{\theta^{*}}_{\mu^{*}_{j}}||_{\infty}+||Q^{\theta^{*}}_{\mu^{*}_{j+1}}-Q^{\theta^{*}}_{\mu_{j+1}}||_{\infty}+\frac{1}{n}\sum_{i=1}^{n}\max_{(s,a)\in\mathcal{S}\times\mathcal{A}}|Q^{\theta^{*}}_{\mu_{j+1}}(s,a)-Q^{{\theta}_{ki}}_{\mu_{j+1}}(s,a)|$
$\leq\gamma||Q^{*}-Q^{\theta^{*}}_{\mu^{*}_{j}}||_{\infty}+||Q^{\theta^{*}}_{\mu^{*}_{j+1}}-Q^{\theta^{*}}_{\mu_{j+1}}||_{\infty}+\frac{1}{n}\sum_{i=1}^{n}|\theta_{ki}-\theta^{*}|\max_{(s,a)\in\mathcal{S}\times\mathcal{A}}K_{\mu_{j+1}}(s,a)$
$\quad[\text{assuming $Q_{\mu}^{\theta}$ to be Lipschitz continuous w.r.t $\theta$, $\exists K_{\mu}<\infty$ for all $\mu$}]$
$=\gamma||Q^{*}-Q^{\theta^{*}}_{\mu^{*}_{j}}||_{\infty}+||Q^{\theta^{*}}_{\mu^{*}_{j+1}}-Q^{\theta^{*}}_{\mu_{j+1}}||_{\infty}+\frac{1}{n}\sum_{i=1}^{n}|\theta_{ki}-\theta^{*}|K^{\prime}_{\mu_{j+1}}$
$\quad[\text{where $K^{\prime}_{\mu}=\max_{(s,a)}K_{\mu(s,a)}$}]$
$=\gamma||Q^{*}-Q^{\bm{\theta^{(k)}}}_{\mu_{j}}+Q^{\bm{\theta^{(k)}}}_{\mu_{j}}-Q^{\theta^{*}}_{\mu_{j}}+Q^{\theta^{*}}_{\mu_{j}}-Q^{\theta^{*}}_{\mu^{*}_{j}}||_{\infty}+||Q^{\theta^{*}}_{\mu^{*}_{j+1}}-Q^{\theta^{*}}_{\mu_{j+1}}||_{\infty}$
$\quad+\frac{1}{n}\sum_{i=1}^{n}|\theta_{ki}-\theta^{*}|K^{\prime}_{\mu_{j+1}}$
$\leq\gamma||Q^{*}-Q^{\bm{\theta^{(k)}}}_{\mu_{j}}||_{\infty}+\gamma||Q^{\theta^{*}}_{\mu^{*}_{j}}-Q^{\theta^{*}}_{\mu_{j}}||_{\infty}+||Q^{\theta^{*}}_{\mu^{*}_{j+1}}-Q^{\theta^{*}}_{\mu_{j+1}}||_{\infty}$
$\quad+\gamma||Q^{\bm{\theta^{(k)}}}_{\mu_{j}}-Q^{\theta^{*}}_{\mu_{j}}||_{\infty}+\frac{1}{n}\sum_{i=1}^{n}|\theta_{ki}-\theta^{*}|K^{\prime}_{\mu_{j+1}}$
$\leq\gamma||Q^{*}-Q^{\bm{\theta^{(k)}}}_{\mu_{j}}||_{\infty}+\gamma||Q^{\theta^{*}}_{\mu^{*}_{j}}-Q^{\theta^{*}}_{\mu_{j}}||_{\infty}+||Q^{\theta^{*}}_{\mu^{*}_{j+1}}-Q^{\theta^{*}}_{\mu_{j+1}}||_{\infty}$
$\quad+\frac{\gamma}{n}\sum_{i=1}^{n}|\theta_{ki}-\theta^{*}|K^{\prime}_{\mu_{j}}+\frac{1}{n}\sum_{i=1}^{n}|\theta_{ki}-\theta^{*}|K^{\prime}_{\mu_{j+1}}$
$=\gamma||Q^{*}-Q^{\bm{\theta^{(k)}}}_{\mu_{j}}||_{\infty}+\gamma||Q^{\theta^{*}}_{\mu^{*}_{j}}-Q^{\theta^{*}}_{\mu_{j}}||_{\infty}+||Q^{\theta^{*}}_{\mu^{*}_{j+1}}-Q^{\theta^{*}}_{\mu_{j+1}}||_{\infty}$
$\quad+\frac{1}{n}\sum_{i=1}^{n}|\theta_{ki}-\theta^{*}|(\gamma K^{\prime}_{\mu_{j}}+K^{\prime}_{\mu_{j+1}})$
$=\gamma||Q^{*}-Q^{\bm{\theta^{(k)}}}_{\mu_{j}}||_{\infty}+\gamma\zeta_{j}(k,n)$
$=\gamma^{j}||Q^{*}-Q^{\bm{\theta^{(k)}}}_{\mu_{1}}||_{\infty}+\sum_{l=1}^{j}\gamma^{j-l+1}\zeta_{l}(k,n)$
where $\zeta_{j}(k,n)=||Q^{\theta^{*}}_{\mu^{*}_{j}}-Q^{\theta^{*}}_{\mu_{j}}||_{\infty}+\frac{1}{\gamma}||Q^{\theta^{*}}_{\mu^{*}_{j+1}}-Q^{\theta^{*}}_{\mu_{j+1}}||_{\infty}+\frac{1}{n}\sum_{i=1}^{n}|\theta_{ki}-\theta^{*}|(K^{\prime}_{\mu_{j}}+\frac{1}{\gamma}K^{\prime}_{\mu_{j+1}})$
Let us define, $Y_{k}=\theta-\hat{\theta_{k}}$, where $\theta\sim\pi_{k}$ and $\hat{\theta_{k}}$ denotes the Scoring rule minimizer obtained after observing $k$th episode. So, according to the Bernstein–von Mises (BvM) theorem, $Y_{k}=O_{p}((k\tau)^{-1/2})$ assuming $Y_{k}$ has finite expectation. Furthermore, from the consistency results of M-estimators (Van der Vaart, 2000), under certain regularity conditions,$\hat{\theta_{k}}$ converges to the expected Scoring Rule minimizer ($\theta^{*}$) at the rate of $O_{p}((k\tau)^{-1/2})$.
For the $n$ samples drawn from the $k$th posterior, if we define $Y_{ki}=|\theta_{ki}-\hat{\theta_{k}}|+|\hat{\theta_{k}}-{\theta^{*}}|$ then, $Y_{ki}=O_{p}((k\tau)^{-1/2})$ for all $i=1,2,\ldots n$. Hence, $1/n\sum_{i=1}^{n}|Y_{ki}|=O_{p}((nk\tau)^{-1/2})$. Therefore, the last term in the equation (A.2) vanishes with large $k$ and $n$. Thus, compared to the classical TS, for ETS the last term shrinks $\sqrt{n}$ times faster.
Note that, $\mu^{*}_{j}(s)=\arg\max_{a}Q_{\mu^{*}_{j-1}}^{\theta^{*}}(s,a)$ and $\mu_{j}(s)=\arg\max_{a}Q_{\mu_{j-1}}^{\bm{\theta(k)}}(s,a)$. So for large enough $k$ and $n$, from the consistency of posterior mean, we can say that, $Q_{\mu^{*}_{j}}^{\theta^{*}}\approx Q_{\mu_{j}}^{\bm{\theta(k)}}$ for any policy iteration step $j$. Then, the second and third term in equation (A.2) also vanish as $k\rightarrow\infty$ and $n\rightarrow\infty$.
A.3 Tuning parameters for posterior sampling using SMC
Parameter Inverted Pendulum Well specified Inverted Pendulum Misspecified Hopper
$\epsilon$(adSGLD step size) $10^{-2}$ $10^{-5}$ $10^{-6}$
$a$ (adSGLD parameter) $10^{-2}$ $10^{-4}$ $10^{-6}$
$\kappa$ (ESS/ CESS multiplier) $0.9$ $0.9$ $0.9$
Number of adSGLD moves per iteration of SMC $10$ $20$ $5$
Number of simulations to estimate the SR loss $10$ $10$ $10$
Table 1: Tuning parameters for SMC
A.4 Zeroth order gradient
Often, simulator models are not differentiable, yet we seek to leverage gradient information for improved sampling. In such cases, a gradient-free optimization technique involving zeroth order (ZO) gradient (Liu et al., 2020) can be used. The multi-point ZO gradient estimate of a function $f(\theta)$ is defined as,
$\widehat{\nabla_{\theta}}U(\theta)=\frac{1}{\mu b}\sum_{i=1}^{b}[U(\theta+\mu z_{i})-f(\theta)]z_{i}$
with approximation error $O(\frac{d}{b})||\nabla_{\theta}U(\theta)||_{2}^{2}+O(\frac{\mu^{2}d^{3}}{b})+O({\mu^{2}d})$ (Berahas et al., 2022); where $\{z_{i}\}_{i=1}^{b}$ denotes $b$ i.i.d. samples drawn from $\mathcal{N}(0,I_{d})$ and $\mu$ is a tuning parameter. The above method produces an unbiased estimate of the gradient of the of the smoothed version of $U$ over a random perturbation $Z\sim\mathcal{N}(0,I_{d})$ with smoothing parameter $\mu$, $U_{\mu}(\theta)=\mathbb{E}_{Z}[U(\theta+\mu Z)]$. Intuitively, the final term in the error can be viewed as the bias of the estimate, which tends to grow with the dimension of $\theta$. Meanwhile, the first two terms stem from the variance of the estimate, which can be controlled by augmenting the sample size $b$. Besides, the gradient estimate gets better as $\mu$ is taken to be small, however, in practice the gradient estimate can be affected by system noise if $\mu$ is too small and so the efficiency of the estimate relies on the careful tuning of the smoothing parameter $\mu$. We set $\mu=0.0001$ and $b=30$ for our implementation.
A.5 Implementation details of REINFORCE
We implement REINFORCE for policy learning by drawing actions from three independent Gaussian distributions. These actions are transformed from $\mathbb{R}$ to $[-1,1]$ space as the action space for ‘Hopper’ is defined as $[-1,1]^{3}$. We train a policy network to predict the mean and standard deviation of the action distributions as a function of the observed state. This network comprises three fully connected layers with 32, 64, and 64 neurons, respectively. The policy parameters are updated using one step of the Adam optimizer (setting the learning rate to be $0.0001$) after observing each episode of interaction data. Also we set the discount factor $\gamma=0.99$ to define the discounted return.
When integrating REINFORCE with ETS, the policy network architecture remains unchanged. After each real interaction episode with the environment, we train a simulator model to predict the next state given a state-action pair. Using this simulator, we simulate $500$ interaction episodes, updating the policy at each simulated episode with classical REINFORCE. The updated policy is then used in the next real interaction episode, after which the simulator model is retrained. This process alternates between model learning and policy updates until the $15$th episode.